### v2

# The $L_{p}$-error rate for randomized quasi-Monte Carlo self-normalized importance sampling of unbounded integrands

Abstract
Self-normalized importance sampling (SNIS) is a fundamental tool in Bayesian inference when the posterior distribution involves an unknown normalizing constant. Although $L_{1}$-error (bias) and $L_{2}$-error (root mean square error) estimates of SNIS are well established for bounded integrands, results for unbounded integrands remain limited, especially under randomized quasi-Monte Carlo (RQMC) sampling. In this work, we derive $L_{p}$-error rate $(p\geq 1)$ for RQMC-based SNIS (RQMC-SNIS) estimators with unbounded integrands on unbounded domains. A key step in our analysis is to first establish the $L_{p}$-error rate for plain RQMC integration. Our results allow for a broader class of transport maps used to generate samples from RQMC points. Under mild function boundary growth conditions, we further establish $L_{p}$-error rate of order $\mathcal{O}(N^{-\beta+\epsilon})$ for RQMC-SNIS estimators, where $\epsilon>0$ is arbitrarily small, $N$ is the sample size, and $\beta\in(0,1]$ depends on the boundary growth rate of the resulting integrand. Numerical experiments validate the theoretical results.
Key words: Quasi-Monte Carlo, Self-normalized importance sampling, $L_{p}$-error rate, unbounded integrand.
1 Introduction
Monte Carlo (MC) method is a standard approach for approximating expectations of the form $\pi(f):=\mathbb{E}_{\pi}[f(\bm{X})]$, where $\pi$ is the target distribution and $f$ is a test function. MC has been extensively applied in areas such as engineering, finance, and statistics [glasserman2004]. Since MC requires sampling from $\pi$, situations where direct sampling is infeasible, most notably in Bayesian posterior inference, are typically addressed using importance sampling (IS). Given a tractable proposal distribution $q$, the standard IS estimator is defined as
${I}_{N}^{\mathrm{IS}}(f)=\frac{1}{N}\sum_{i=1}^{N}\omega(\bm{x}_{i})f(\bm{x}_{i}),$ (1)
where $N$ is the sample size, $\bm{x}_{i}\sim q$, and $\omega(\bm{x}_{i})=\pi(\bm{x}_{i})/{q(\bm{x}_{i})}$ is the weights. However, in most practical Bayesian problems, the posterior density $\pi(\bm{x})$ is known only up to a normalizing constant, rendering the standard IS estimator infeasible. This issue can be addressed by the self-normalized importance sampling (SNIS) estimator, which normalizes the weights:
$\pi_{N}(f)=\sum_{i=1}^{N}\frac{\omega(\bm{x}_{i})}{{\sum_{j=1}^{N}\omega(\bm{x}_{j})}}f(\bm{x}_{i}).$ (2)
For bounded test functions, root mean square error (RMSE) bounds under independent and identically distributed (IID) sampling are well established [sanz2023]. Agapiou et al. [agapiou2017] extended these results by deriving RMSE bounds for unbounded test functions. Deligiannidis et al. [deligiannidis2024] further generalized the analysis to arbitrary $L_{p}$-errors $(p\geq 2)$ under suitable moment conditions on both the test function $f$ and the weight function $\omega$. In all cases, the RMSE converges at rate $\mathcal{O}(N^{-1/2})$, and the $L_{p}$-error established in [deligiannidis2024] also achieves this rate, which may be slow for practical applications.
Quasi-Monte Carlo (QMC) method is a deterministic version of MC, known for its superior convergence rate compared to standard MC [bookdick2010]. Instead of using IID random sequences, QMC uses deterministic low-discrepancy sequences. For a $d$-dimensional integral, QMC achieves deterministic error bounds of order $\mathcal{O}(N^{-1}(\log N)^{d})$ for certain regular functions, which is asymptotically superior to the MC rate of $\mathcal{O}(N^{-1/2})$. In practice, randomized QMC (RQMC) is more commonly used. It preserves the low discrepancy property of QMC and allows error estimation via multiple replications. QMC and RQMC methods have been successfully applied in finance and statistics [he2023, le2009, wang2017]. Given the advantages of RQMC over MC, several works have studied the incorporation of IS and SNIS with RQMC methods. For RQMC-based IS (RQMC-IS), Zhang et al. [Zhang2021] investigated effective strategies for combining IS with conditional Monte Carlo and dimension-reduction techniques within the RQMC framework. Under the “QMC-friendly” boundary growth conditions, He et al. [he2023] proved the RMSE of RQMC-IS achieves a convergence rate of $\mathcal{O}(N^{-1+\epsilon})$, where $\epsilon>0$ is arbitrarily small, for unbounded functions defined on unbounded domains. The work [he2023] also investigated the effect of different proposal distributions on the performance of RQMC-IS. The projection-based framework proposed by Ouyang et al. [ouyang2024] provides an alternative approach for analyzing the RMSE convergence rate of RQMC-IS.
However, existing results for RQMC-based SNIS (RQMC-SNIS) do not cover the case of unbounded test functions on unbounded domains. Dick et al. [dick2013] derived an explicit error bound for RQMC-SNIS estimator using a weighted star-discrepancy. But their result requires sufficiently regular conditions on the functions $\omega$ and $f$, which are not satisfied when $f$ is unbounded. Gerber and Chopin [sqmc2015] demonstrated the validity of RQMC-SNIS for unbounded test functions on bounded domains by introducing an extreme norm on $0,1)^{d}$. However, extending this approach to unbounded domains is challenging. He et al. [he2024] established an RMSE estimate for RQMC-SNIS under the assumption that the estimator $\pi_{N}(f)$ is uniformly bounded. To the best of our knowledge, for unbounded test functions on unbounded domains, even when the test functions are sufficiently smooth, there are currently no results on the $L_{1}$-error rate for the RQMC-SNIS estimator, let alone on the $L_{p}$-error rate for $p\geq 2$ . In this work, we study the $L_{p}$-error rate for the RQMC-SNIS estimator allowing for unbounded test functions on unbounded domains.
A key step in establishing $L_{p}$-error rate for MC-based SNIS [deligiannidis2024] is to estimate the $L_{p}$-errors of $I_{N}^{\mathrm{IS}}(f)$ and $I_{N}^{\mathrm{IS}}(1)$ defined in (1). For IID sampling, these estimates are relatively straightforward and can be derived using the Marcinkiewicz–Zygmund inequality [ferger2014]. In contrast, for RQMC, where $\bm{x}_{i}=\mathcal{T}(\bm{u}_{i})\sim q$ with RQMC points $\{\bm{u}_{i}\}_{i=1}^{N}$ and a transport map $\mathcal{T}:(0,1)^{d}\to\mathbb{R}^{d}$, establishing analogous $L_{p}$-error rate is substantially more challenging. Existing results on the $L_{p}$-error rate for RQMC are limited to the inversion transformation $\mathcal{T}=\Phi^{-1}$, where $\Phi$ denotes the cumulative distribution function (CDF) of the standard normal distribution, and to test functions with sub-quadratic growth [chen2025]. Our first contribution is to extend these results to a much broader class of transport maps, including shifted, scaled, and commonly used nonlinear transformations, as well as to a wider class of test functions satisfying the critical quadratic growth. Moreover, establishing $L_{p}$-error rate for RQMC with general transport maps and a broader class of test functions is crucial for supporting analytical tools such as higher-order Markov and Hölder inequalities, as well as the estimation of skewness and kurtosis in uncertainty quantification.
Based on the established $L_{p}$-error rate for plain RQMC integration, our second contribution derives the $L_{p}$-error rate for RQMC-SNIS under suitable moment conditions. The resulting rate is of order $\mathcal{O}(N^{-\beta+\epsilon})$, where $\beta\in(0,1]$ depends on the growth rate of $\omega f$. For the function $\omega f$ belonging to a broad class of “QMC-friendly”, $\beta$ approaches $1$, yielding a nearly $\mathcal{O}(N^{-1})$ convergence rate. To the best of our knowledge, this is the first systematic $L_{p}$-error analysis of RQMC-SNIS for unbounded test functions on unbounded domains. Setting $p=1$ yields the bias rate, while $p=2$ corresponds to the RMSE rate. This contribution bridges a key theoretical gap in the analysis of RQMC-SNIS and substantially broadens its applicability to practical Bayesian computation.
The paper is organized as follows. Section 2 provides a brief introduction to QMC and RQMC. Section 3 derives the $L_{p}$-error rate for RQMC. Section 4 extends the analysis to RQMC-SNIS. Section 5 reports numerical experiments that validate the theoretical findings. Conclusions are drawn in Section 6. Some applications of the $L_{p}$-error rate can be found in Appendix A.
2 Background and notation
2.1 Notation
In this work, all norms are taken to be Euclidean norms. To prevent confusion, we adopt the convention that bold symbols (e.g., $\bm{u}$, $\bm{v}$) denote vectors, while ordinary symbols denote scalars. For instance, $\bm{u}_{j}$ refers to the $j$-th point in a low-discrepancy point set, whereas $u_{j}$ denotes the $j$-th component of a $d$-dimensional vector $\bm{u}$. Denote $1\!:\!d={1,2,\ldots,d}$. Throughout, $\bm{v}$ will always denote a subset of 1:d1\!:\!d, $\bm{u}$ a point in $(0,1)^{d}$, and $\bm{x}$ a point in $\mathbb{R}^{d}$. The notation $\bm{a}_{\bm{v}}:\bm{b}_{\bar{\bm{v}}}$ denotes a vector in $\mathbb{R}^{d}$ whose $j$-th component is $a_{j}$ if $j\in\bm{v}$ and $b_{j}$ otherwise. And, $\partial^{\bm{v}}h(\bm{u})$ denotes the mixed partial derivative of $h(\bm{u})$ with respect to $u_{j}$ with $j\in\bm{v}$. For $\bm{a}=\left(a_{1},\ldots,a_{d}\right)\in\mathbb{N}_{0}^{d}$ where $\mathbb{N}_{0}^{d}$ is the set of all $d$-dimensional vectors of non-negative integers, define $|\bm{a}|=a_{1}+\cdots+a_{d}$, and
$D^{\bm{a}}f(\bm{x}):=\frac{\partial^{|\bm{a}|}}{\partial x_{1}^{a_{1}}\ldots\partial x_{d}^{a_{d}}}f(\bm{x}),\quad\forall\bm{x}\in\mathbb{R}^{d}.$
Moreover, we denote by $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ the smallest and largest eigenvalues of a matrix, respectively. For functions $f$ and $g$, we write $f\lesssim g$ if there exists a constant $c>0$ such that $f\leq cg$. And $f=\mathcal{O}(g)$ means there exists a constant $c>0$ such that $|f|\leq c|g|$. In this case, we also write $g=\Omega(f)$.
2.2 Background
In this section, we provide a brief introduction to (R)QMC and review some recent results on the convergence rates of RQMC method. Further details can be found in the monographs [bookdick2010, booknieder1992].
QMC is commonly used for the numerical evaluation of $d$-dimensional integral $\int_{(0,1)^{d}}f({\color[rgb]{0,0,0}\bm{u}})\mathrm{d}{\color[rgb]{0,0,0}\bm{u}}$. The estimator is $\frac{1}{N}\sum_{i=1}^{N}f(\bm{u}_{i})$, where $\bm{u}_{i}\in(0,1)^{d}$. In comparison to MC, where $\{\bm{u}_{i}\}_{i=1}^{N}$ are IID points, $\{\bm{u}_{i}\}_{i=1}^{N}$ used in QMC are low-discrepancy points. The discrepancy of a point set is typically quantified using the star discrepancy, which is defined as
$D_{N}^{*}\left(\bm{u}_{1},\ldots,\bm{u}_{N}\right)=\sup_{\bm{z}\in(0,1)^{d}}\left|\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\left\{\bm{u}_{i}\in[\mathbf{0},{\color[rgb]{0,0,0}\bm{z}})\right\}}-\prod_{j=1}^{d}z_{j}\right|.$ (3)
A point set with a star discrepancy of $\mathcal{O}(N^{-1}(\log N)^{d-1})$ is referred to as a low-discrepancy point set. Commonly used low-discrepancy point sets include Sobol’ point sets, Faure point sets, and others. The classical Koksma-Hlawka (K-H) inequality provides an error bound on the integration error:
$\left|\int_{(0,1)^{d}}f({\color[rgb]{0,0,0}\bm{u}})\mathrm{d}{\color[rgb]{0,0,0}\bm{u}}-\frac{1}{N}\sum_{i=1}^{N}f(\bm{u}_{i})\right|\leq D_{N}^{*}\left(\bm{u}_{1},\ldots,\bm{u}_{N}\right)V_{\mathrm{HK}}(f),$ (4)
where $V_{\mathrm{HK}}(f)$ denotes the variation of $f$ in the sense of Hardy and Krause. For its definition and properties, we refer to [owen2005hk]. A function $f$ is said to be of bounded variation in the sense of Hardy and Krause ($\mathrm{BVHK}$) if $V_{\mathrm{HK}}(f)$ is finite. To facilitate error estimation, RQMC is commonly used. Typical randomization methods include random shifts [cranley1976], digital shifts [owenqmc] and scrambling [owen1995].
To derive the convergence rate of RQMC method, Owen [owen2006] introduced a growth condition for unbounded function defined on $(0,1)^{d}$ and established the $L_{1}$-error rate using a low-variation extension strategy. Building on this approach, He et al. [he2023] further derived the $L_{2}$ convergence rate by exploiting the fact that the scrambled net variance is no worse than a constant times the MC variance.
Recently, Ouyang et al. [ouyang2024] refined the growth condition for unbounded functions on $\mathbb{R}^{d}$, focusing on the standard normal target $\pi\sim\mathcal{N}(\bm{0},I_{d})$, and established the RMSE convergence rates for RQMC using a projection operator. Chen et al. [chen2025] showed that the projection method can also be employed to derive the $L_{p}$-error rate, thereby extending the error analysis of RQMC with $\pi\sim\mathcal{N}(\bm{0},I_{d})$.
The problem considered in this work is to establish the $L_{p}$-error rate of the RQMC-SNIS estimator, given by:
$\left[\mathbb{E}\left(\left|\pi_{N}(f)-\pi(f)\right|^{p}\right)\right]^{1/p},$
where $\pi(f)=\mathbb{E}_{\pi}[f(\bm{X})]$, $\pi_{N}(f)$ is the SNIS estimator defined in (2) with $\bm{x}_{i}=\mathcal{T}(\bm{u}_{i})\sim q,~i=1,2,\ldots,N$. Here, $\{\bm{u}_{i}\}_{i=1}^{N}$ is an RQMC point set and $\mathcal{T}:(0,1)^{d}\to\mathbb{R}^{d}$ is a transport map. The key step in our analysis is estimating the $L_{p}$-error rate for plain RQMC integration. In the following section, we build upon the projection method proposed by [ouyang2024] to establish the $L_{p}$-error rate for RQMC under a general class of functions and transport maps.
3 The $L_{p}$-error rate of RQMC
In the section, we consider the $L_{p}$-error rate for RQMC integration as follows:
$\left[\mathbb{E}\left(\left|I_{N}(f)-\pi(f)\right|^{p}\right)\right]^{1/p},$
where
$I_{N}(f)=\frac{1}{N}\sum_{i=1}^{N}f(\bm{x}_{i})=\frac{1}{N}\sum_{i=1}^{N}f(\mathcal{T}(\bm{u}_{i})),$ (5)
and $\{\bm{u}_{1},\ldots,\bm{u}_{N}\}$ is an RQMC point set satisfying
$\mathbb{P}\left(D_{N}^{*}(\bm{u}_{1},\ldots,\bm{u}_{N})=\mathcal{O}(N^{-1}(\log N)^{(d-1)})\right)=1,$ (6)
and each $\bm{u}_{i}\sim\mathcal{U}(0,1)^{d}$, $\bm{x}_{i}=\mathcal{T}(\bm{u}_{i})\sim\pi$. We firstly focus on the function $f$ satisfying the following Assumption 1 with $M\in\mathbb{R}$, and the inversion transformation as the transport map
$\mathcal{T}(\bm{u})=G^{-1}(\bm{u})=\left(G_{1}^{-1}(u_{1}),G_{2}^{-1}(u_{2}),\ldots,G_{d}^{-1}(u_{d})\right),$ (7)
where $G_{i}$ denotes the CDF of a uni-variate distribution with density $g_{i}$, $i=1,2,\ldots,d$. General transport maps will be considered subsequently.
Assumption 1.
Suppose that $f(\bm{x})$ is a real-valued continuous function on $\mathbb{R}^{d}$ and $D^{\bm{a}}f$ exists for any $\bm{a}\in\mathbb{N}_{0}^{d}$ satisfying $|\bm{a}|\leq d$. Assume that there exist a growth rate $M\in\mathbb{R}$ such that, for any $\bm{a}\in\mathbb{N}_{0}^{d}$ satisfying $|\bm{a}|\leq d$,
$\left|D^{\bm{a}}f(\bm{x})\right|\lesssim\exp\left\{M\|\bm{x}\|^{2}\right\},\quad\forall\bm{x}\in\mathbb{R}^{d}.$
Remark 1.
Ouyang et al. [ouyang2024] considered a smaller function class with sub-quadratic growth $0<k<2$, assuming $|\partial^{\bm{v}}f(x)|\lesssim\exp\{M\|x\|^{k}\}$ for any $\bm{v}\subseteq 1:d$. In contrast, the test function $f$ used in this work satisfies Assumption 1 with the critical exponent $k=2$, which extends the previous setting. This enlarged function space enables the analysis of a wider range of integrands, such as the exponential quadratic functions commonly arising in importance sampling with Gaussian proposals.
If $M\leq 0$, the function $f$ is obviously bounded. To control an unbounded function $f$ with growth rate $M>0$, we require the transport map $\mathcal{T}$ satisfying the following assumption.
Assumption 2.
Let $\bm{X}=\mathcal{T}(\bm{u})$, where $\bm{u}\sim\mathcal{U}(0,1)^{d}$. There exists constants $t_{0}$, $\eta\in\mathbb{R}$ and $\alpha>0$ such that $\mathbb{P}\left(\|\bm{X}\|>t\right)\lesssim t^{\eta}\exp(-\alpha t^{2})$ for any $t>t_{0}$.
Remark 2.
Assumption 2 is essentially a sub-Gaussian-type tail condition [vershynin2018]. It ensures that the probability of $\|\bm{X}\|$ being large decays rapidly enough, which prevents the expectation from blowing up.
Here, we recall Lemma A.2 from [ouyang2024] for the use of following proofs.
Lemma 3.1.
For any positive integer $s$, constants $a>0$ and $t>1/\sqrt{2a}$, we have
$\int_{t}^{\infty}x^{s}\exp\left\{-ax^{2}\right\}dx\leq\frac{(s+2)!!}{4a}t^{d-1}\exp\left\{-at^{2}\right\}.$
Proof.
See the proof of Lemma A.2 in [ouyang2024]. ∎
Remark 3.
When $\bm{X}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$, then its density $g(\bm{x})\lesssim\exp(-\frac{\lambda}{2}\|\bm{x}\|^{2})$ with $\lambda=\lambda_{\min}(\bm{\Sigma}^{-1})-\varepsilon$ and any $\varepsilon>0$. By the spherical coordinate transformation, for $t>1/\sqrt{\lambda}$, we have
$\mathbb{P}\left(\|\bm{X}\|>t\right)=\int_{\|\bm{x}\|>t}g(\bm{x})d\bm{x}\lesssim\int_{t}^{\infty}\exp\left(-\frac{\lambda}{2}r^{2}\right)r^{d-1}dr\lesssim t^{d-2}\exp\left(-\frac{\lambda}{2}t^{2}\right).$ (8)
In the last inequality, Lemma 3.1 was used. Hence, a Gaussian random variable satisfies Assumption 2 with $\eta=d-2$, $\alpha=\lambda/2$, $t_{0}=1/\sqrt{\lambda}$.
We now revisit the projection method proposed in [ouyang2024]. The main idea behind the RMSE convergence result in [ouyang2024] is the use of a smoothed projection operator ${P}_{r}$, where $r>1$ denotes the projection radius. Following [ouyang2024], in the one-dimensional setting, the operator ${P}_{r}:\mathbb{R}\rightarrow\mathbb{R}$ is defined piecewise as follows:
$P_{r}(x)=\begin{cases}-r+\frac{1}{2},&x\in(-\infty,-r],\\ \frac{1}{2}x^{2}+rx+\frac{(r-1)^{2}}{2},&x\in(-r,-r+1),\\ x,&x\in[-r+1,r-1],\\ -\frac{1}{2}x^{2}+rx-\frac{(r-1)^{2}}{2},&x\in(r-1,r),\\ r-\frac{1}{2},&x\in[r,\infty).\end{cases}$ (9)
For the multivariate case, the operator acts component-wise as
${P}_{r}(\bm{x})=({P}_{r}(x_{1}),\ldots,{P}_{r}(x_{d})).$
Using this operator, Ouyang et al. [ouyang2024] establishes the $L_{2}$-error rate for RQMC integration with $\mathcal{T}=\Phi^{-1}$.
The original projection operator ${P}_{r}$, defined component-wise, leads to $\|{P}_{r}(\bm{x})\|\leq\sqrt{d}r$ for any $\bm{x}\in\mathbb{R}^{d}$, which will introduce an undesirable dimension factor in the subsequent error order. It is worthy noting that, to simultaneously ensure that the upper bound of the operator norm is dimension-independent and that the range of $\|\bm{x}\|$ over which the projection operator coincides with the identity map is also independent of the dimension, a component-wise definition for the projection operator is no longer feasible. Therefore, we construct a novel radially defined $d$-dimensional projection operator that satisfies both properties.
To ensure a smooth transition, we firstly focus on the following function,
$\psi(t)=\frac{\int_{0}^{t}s^{d}(1-s)^{d}ds}{\int_{0}^{1}s^{d}(1-s)^{d}ds},~0\leq t\leq 1,$ (10)
which is the CDF of a symmetric Beta distribution with parameter $d+1$. It is straightforward to verify that $0\leq\psi(t)\leq 1$, $\psi(0)=0,\psi(1)=1$.
Lemma 3.2.
For the function $\psi$ defined in (10), $D^{k}\psi(0)=D^{k}\psi(1)=0,1\leq k\leq d.$ Moreover, for $0<t<1$, there exists a constant $C>0$ such that $|D^{k}\psi(t)|\leq C,1\leq k\leq d$.
Proof.
Denote $\zeta(s)=s^{d}(1-s)^{d}$ and $c=1/\int_{0}^{1}\zeta(s)ds=\frac{(2d+1)!}{(d!)^{2}}$, so that $\psi(t)=c\int_{0}^{t}\zeta(s)ds$. By the Leibnitz formula, $D^{k}\psi(t)=cD^{k-1}\zeta(t),1\leq k\leq d$, where
$D^{k-1}\zeta(t)$ $=\sum_{i=0}^{k-1}\binom{k-1}{i}D^{i}(t^{d})D^{k-1-i}((1-t)^{d}).$
Evaluating at the boundaries, for $1\leq k\leq d$, we have
$D^{k}\psi(0)=cD^{k-1}\zeta(0)=0,\quad D^{k}\psi(1)=cD^{k-1}\zeta(1)=0.$
Moreover, for $0\leq t\leq 1$, each term $D^{i}(t^{d}),D^{k-1-i}((1-t)^{d})$ is bounded by a constant depending only on $d$ and $k$. Hence, for $1\leq k\leq d$, there exists a constant $C>0$ such that
$|D^{k}\psi(t)|=|cD^{k-1}\zeta(t)|\leq C.$
This completes the proof. ∎
Then, our radially defined $d$-dimensional projection operator is defined as
$P_{r}(\bm{x})=\begin{cases}\bm{x},&\|\bm{x}\|\leq(1-\delta)r,\\ (1-\psi(t(\bm{x})))\bm{x},&(1-\delta)r<\|\bm{x}\|<r,\\ \bm{0},&\|\bm{x}\|\geq r,\end{cases}$ (11)
where $(1-\delta)r>1$, $t(\bm{x})=(\|\bm{x}\|-(1-\delta)r)/(\delta r)$, and $0<\delta<1/2$ is a sufficiently small constant.
Remark 4.
A comparison of the two projection operators is given in Figure 1, where we illustrate the first component $P_{r}^{(1)}(\bm{x})$ of the two-dimensional projection operator with $d=2,r=10,\delta=0.1$. It can be observed that, in our operator (11), the function value of the operator depends on both components, whereas the operator of [ouyang2024] depends only on the first component.
(a) Our projection operator (11)
(b) The projection operator (9) proposed by [ouyang2024]
Figure 1: A comparison on the first components of the two projection operators in $\mathbb{R}^{2}$ with $r=10$ for both operators and $\delta=0.1$ for our proposed operator.
Lemma 3.3.
For any $\varnothing\neq\bm{v}\subseteq 1:d$ and $(1-\delta)r<\|\bm{x}\|<r$, $|\partial^{\bm{v}}\psi(t(\bm{x}))|\lesssim r^{-1}$.
Proof.
By Faa di Bruno formula [cons1996], $\partial^{\bm{v}}\psi(t(\bm{x}))$ is a sum of finite terms of the form
$\psi^{(m)}(t)\prod_{l=1}^{m}\partial^{\bm{\ell}_{l}}t(\bm{x}),$
where $1\leq m\leq|\bm{v}|$, and $\bm{\ell}_{l}$ are nonempty disjoint sets satisfying $\cup_{l=1}^{|m|}\bm{\ell}_{l}=\bm{v}$. By the definition of $t(\bm{x})$, we have
$\partial^{\bm{\ell}}t(\bm{x})=\partial^{\bm{\ell}}\left[\frac{\|\bm{x}\|-(1-\delta)r}{\delta r}\right]=\frac{\partial^{\bm{\ell}}(\|\bm{x}\|)}{\delta r}=\frac{1}{\delta r}\frac{P_{|\bm{\ell}|}(\bm{x})}{\|\bm{x}\|^{2|\bm{\ell}|-1}}\leq\frac{M_{|\bm{\ell}|}}{\delta r}\frac{1}{\|\bm{x}\|^{|\bm{\ell}|-1}},$ (12)
[…]
$|P_{l}(\bm{x})|=\|\bm{x}\|^{l}P_{l}\left(\frac{\bm{x}}{\|\bm{x}\|}\right).$
Since the unit sphere is compact, $P_{l}$ attains its maximum on the unit sphere. There exists a constant $M_{l}$ such that $|P_{l}(\bm{x})|\leq M_{l}\|\bm{x}\|^{l}$. By Lemma 3.2, it follows that
$\left|\psi^{(m)}(t)\prod_{l=1}^{m}\partial^{\bm{\ell}_{l}}t(\bm{x})\right|\leq C\prod_{l=1}^{m}\frac{M_{|\bm{\ell}_{l}|}}{\delta r}\frac{1}{\|\bm{x}\|^{|\bm{\ell}_{l}|-1}}\lesssim\frac{1}{(\delta r)^{m}}\lesssim\frac{1}{r},$
where we used the fact that $\|\bm{x}\|>(1-\delta)r>1$. This completes the proof. ∎
Lemma 3.4.
Let $P_{r}^{(i)}(\bm{x})$ be the $i$-th component of $P_{r}(\bm{x})$, $i=1,2,\ldots,d$. For any $\varnothing\neq\bm{v}\subseteq 1:d$ and $1\leq i\leq d$, the projection $P_{r}$ defined in (11) satisfies:
  1. (i)
    $\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})$ exists for any $\bm{x}\in\mathbb{R}^{d}$,
  2. (ii)
    $\left|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})\right|\lesssim\mathds{1}_{\{\|\bm{x}\|<r\}}$,
  3. (iii)
    For any $\bm{x}\in\mathbb{R}^{d}$, $\|P_{r}(\bm{x})\|\leq\|\bm{x}\|$ and $\|P_{r}(\bm{x})\|\leq(1-\delta/2)r$.
Proof.
We prove the items in order.
(i) In fact, the projection operator can be interpreted as a vector-valued function whose components are scaled by radial factors. Specifically, $P_{r}^{(i)}(\bm{x})=\xi(\rho)x_{i}$ where $\rho=\|\bm{x}\|$,
$\xi(\rho)=\begin{cases}1,&\rho\leq(1-\delta)r,\\[4.0pt] \kappa(\rho),&(1-\delta)r<\rho<r,\\[4.0pt] 0,&\rho\geq r,\end{cases}\quad\text{and}\quad\kappa(\rho)=1-\psi\left(\dfrac{\rho-(1-\delta)r}{\delta r}\right).$
It is straightforward to verify that the smoothness of the projection operator is completely determined by the smoothness of $\xi(\rho)$. By Lemma 3.2, we check that $\xi(\rho)$ and its derivatives up to order $d$ match continuously at $\rho=(1-\delta)r$ and $\rho=r$. Hence, for any $i=1,2,\ldots,d$ and $\varnothing\neq\bm{v}\subseteq 1:d$ , $\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})$ exists for any $\bm{x}\in\mathbb{R}^{d}$.
(ii) we begin by considering the region $(1-\delta)r<\rho<r$ and $|\bm{v}|>1$. Since $P_{r}^{(i)}(\bm{x})=\kappa(\rho)x_{i}$, we have
$\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})$ $=\partial^{\bm{v}}\kappa(\rho)x_{i}+\mathds{1}_{\{i\in\bm{v}\}}\partial^{\bm{v}-\{{i}\}}\kappa(\rho).$
Then, by Lemma 3.3, it follows that
$\left|\partial^{\bm{v}}\kappa(\rho)\right|=$ $\left|\partial^{\bm{v}}\left(1-\psi(t(\bm{x}))\right)\right|=\left|-\partial^{\bm{v}}\left(\psi(t(\bm{x}))\right)\right|\lesssim\frac{1}{r}.$
Therefore,
$|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})|$ $\lesssim\frac{1}{r}|x_{i}|+\frac{1}{r}\lesssim 1,$
where we used the fact that $|x_{i}|\leq\rho\leq r$. Similarly, if $|\bm{v}|=1$,
$|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})|=|\partial^{\bm{v}}\kappa(\rho)|\times|x_{i}|+\mathds{1}_{\{\bm{v}=\{{i}\}\}}|\kappa(\rho)|\lesssim\frac{1}{r}|x_{i}|+\left|1-\psi(t)\right|\lesssim 1.$
where we used the fact that $0\leq\psi(t)\leq 1$. For $\rho\leq(1-\delta)r$ and $\rho\geq r$, it is clear that $\left|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})\right|\leq\mathds{1}_{\{\rho\leq(1-\delta)r\}},$ proving the second part of the theorem.
(iii) It is straightforward to check that for $\rho\leq(1-\delta)r$ and $\rho\geq r$, $\|P_{r}(\bm{x})\|\leq(1-\delta/2)r$. We focus on $\rho\in((1-\delta)r,r)$. Note that $t:=t(\bm{x})=\frac{\rho-(1-\delta)r}{\delta r}$, for $0\leq t\leq 1/2$,
$\|P_{r}(\bm{x})\|=(1-\psi(t))\rho\leq\rho=\delta rt+(1-\delta)r\leq(1-\delta/2)r.$
In addition, note that $\zeta(s)=s^{d}(1-s)^{d}$ is symmetric about $s=1/2$, we have
$\int_{0}^{1/2}\zeta(s)ds=\int_{1/2}^{1}\zeta(s)ds=\frac{1}{2}\int_{0}^{1}\zeta(s)ds.$
Hence, for $1/2<t\leq 1$,
$\psi(t)=\frac{\int_{0}^{t}s^{d}(1-s)^{d}ds}{\int_{0}^{1}s^{d}(1-s)^{d}ds}>\frac{\int_{0}^{1/2}s^{d}(1-s)^{d}ds}{\int_{0}^{1}s^{d}(1-s)^{d}ds}=\frac{1}{2}.$
Therefore,
$\|P_{r}(\bm{x})\|=(1-\psi(t))\rho<\rho/2\leq r/2<(1-\delta/2)r,$
completing the proof. ∎
Using the new projection operator (11), for the transport map $\mathcal{T}$ satisfying (7) and $\{\bm{u}_{1},\ldots,\bm{u}_{N}\}$ satisfying (6), one of our main results is following.
Theorem 1.
Let $p\geq 1$. Assume that $f$ satisfies Assumption 1 with a growth rate $M\in\mathbb{R}$.
  1. (i)
    If $M<0$, then for any $\epsilon>0$,
    $\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right).$
  2. (ii)
    Suppose that $\mathcal{T}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$, and that $0\leq M<\alpha/p$. Then for any $\epsilon>0$,
    $\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\gamma+\epsilon}\right),$
    where $\gamma=1-pM/\alpha.$ If $0\leq M<\alpha/(2p)$, then $\gamma>1/2$.
The factor $N^{\epsilon}$ in Theorem 1 arises from the term $(\log N)^{\max\{\eta/(2p),(3d/2-1)\}}$, which can be absorbed into $N^{\epsilon}$ for arbitrarily small $\epsilon>0$.
We now prove several supporting lemmas needed for the proof of part (ii) of Theorem 1. The proof of (i) is analogous and simpler. By triangle inequality, we first bound the error $\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}$ by three terms,
$\mathbb{E}\big|{I}_{N}(f)-{I}_{N}(f\circ P_{r})\big|^{p},\quad\mathbb{E}\big|{I}_{N}(f\circ P_{r})-\mathbb{E}\left[f\circ P_{r}(\bm{X})\right]\big|^{p}\quad\text{and}\quad{\big|\mathbb{E}\left[f\circ P_{r}(\bm{X})\right]-\pi(f)\big|}^{p}.$
Let $Y_{i}=f(G^{-1}(\bm{u}_{i}))-f(P_{r}(G^{-1}(\bm{u}_{i}))),i=1,2,\ldots,N$. Note that each $\bm{u}_{i}\sim\mathcal{U}(0,1)^{d}$, we have
$\mathbb{E}\big|{I}_{N}(f)-{I}_{N}(f\circ P_{r})\big|^{p}=\mathbb{E}\left|\frac{1}{N}\sum_{i=1}^{N}Y_{i}\right|^{p}\leq\mathbb{E}\left(\frac{1}{N}\sum_{i=1}^{N}\big|Y_{i}\big|^{p}\right)=\mathbb{E}\left|f(\bm{X})-f\circ P_{r}(\bm{X})\right|^{p}.$
By Jensen’s inequality, it follows that
${\big|\mathbb{E}\left[f\circ P_{r}(\bm{X})\right]-\pi(f)\big|}^{p}\leq\mathbb{E}\left|f(\bm{X})-f\circ P_{r}(\bm{X})\right|^{p}.$
So it suffices to examine $\mathbb{E}\left|f(\bm{X})-f\circ P_{r}(\bm{X})\right|^{p}$ to bound the first and third terms.
Lemma 3.5.
Given $p\geq 1$, assume that $\mathcal{T}$ satisfies Assumption 2 with $\alpha>0$, $\eta\in\mathbb{R}$ and $t_{0}=(1-\delta)r>1/\sqrt{2pA}$ and that $f$ satisfies Assumption 1 with $0\leq M<\alpha/p$. Then we have
$\mathbb{E}\left|f\circ P_{r}(\bm{X})-f(\bm{X})\right|^{p}\lesssim r^{\eta}\exp\left(-pA(1-\delta)^{2}r^{2}\right),$ (13)
where $A=\alpha/p-M>0$.
Proof.
By (i) of Lemma 3.4, we have
$\mathbb{E}\left|f\circ P_{r}(\bm{X})-f(\bm{X})\right|^{p}$ $\leq\mathbb{E}\left[\left(\left|f(\bm{\bm{X}})\right|+\left|f\left(P_{r}(\bm{X}\right)\right|\right)^{p}\mathds{1}_{\{\|\bm{X}\|>(1-\delta)r\}}\right]$
$\lesssim\mathbb{E}\left[\exp\left(pM\|\bm{X}\|^{2}\right)\mathds{1}_{\{\|\bm{X}\|>(1-\delta)r\}}\right],$
where we used Assumption 1 with $\bm{a}=\bm{0}$ and the fact that $\|P_{r}(\bm{x})\|\leq\|\bm{x}\|$ by Lemma 3.4. Considering $M>0$ and denoting $Y=\exp\left(pM\|\bm{X}\|^{2}\right)$ and $y_{0}=\exp\left(pM(1-\delta)^{2}r^{2}\right)$, we have
$\mathbb{E}\left[\exp\left(pM\|\bm{X}\|^{2}\right)\mathds{1}_{\{\|\bm{X}\|>(1-\delta)r\}}\right]$
$=$ $~\mathbb{E}\left[Y\mathds{1}_{\{Y>y_{0}\}}\right]=\int_{0}^{\infty}\mathbb{P}\left(Y\mathds{1}_{\{Y>y_{0}\}}>y\right)dy$
$=$ $~\int_{0}^{y_{0}}\mathbb{P}\left(Y\mathds{1}_{\{Y>y_{0}\}}>y\right)dy+\int_{y_{0}}^{\infty}\mathbb{P}\left(Y\mathds{1}_{\{Y>y_{0}\}}>y\right)dy$
$=$ $~y_{0}\mathbb{P}\left(Y>y_{0}\right)+\int_{y_{0}}^{\infty}\mathbb{P}\left(Y>y\right)dy.$
For the first term, by Assumption 2, we have
$y_{0}\mathbb{P}\left(Y>y_{0}\right)$ $=\exp\left(pM(1-\delta)^{2}r^{2}\right)\mathbb{P}\left(\|\bm{X}\|>(1-\delta)r\right)$
$\lesssim r^{\eta}\exp\left(-(\alpha-pM)(1-\delta)^{2}r^{2}\right).$
For the second term, by Assumption 2 again, it follows that
$\int_{y_{0}}^{\infty}\mathbb{P}\left(Y>y\right)dy$ $=\int_{y_{0}}^{\infty}\mathbb{P}\left(\|\bm{X}\|>\left(\frac{\log y}{pM}\right)^{1/2}\right)dy$
$\lesssim\int_{y_{0}}^{\infty}\left(\frac{\log y}{pM}\right)^{\eta/2}\exp\left(-\alpha\left(\frac{\log y}{pM}\right)\right)dy$
$\lesssim\int_{\sqrt{pM}(1-\delta)r}^{\infty}z^{\eta+1}\exp\left(-\frac{\alpha-pM}{pM}z^{2}\right)dz$
$\lesssim r^{\eta}\exp\left(-(\alpha-pM)(1-\delta)^{2}r^{2}\right).$
In the last inequality, Lemma 3.1 was used. For $M=0$, we also have
$\mathbb{E}\left[\exp\left(pM\|\bm{X}\|^{2}\right)\mathds{1}_{\{\|\bm{X}\|>(1-\delta)r\}}\right]=\mathbb{P}\left(\|\bm{X}\|>(1-\delta)r\right)\lesssim r^{\eta}\exp\left(-\alpha(1-\delta)^{2}r^{2}\right).$
Therefore,
$\mathbb{E}\left[\left(f\circ P_{r}(\bm{X})-f(\bm{X})\right)^{p}\right]\lesssim r^{\eta}\exp\left(-pA(1-\delta)^{2}r^{2}\right).$
This completes the proof. ∎
Now, we bound the term $\mathbb{E}\left|{I}_{N}(f\circ P_{r})-\mathbb{E}[f\circ P_{r}(\bm{X})]\right|^{p}$.
Lemma 3.6.
Assume that $f$ satisfies Assumption 1 with $M\geq 0$. Let $h(\bm{u})=(f\circ P_{r})\circ G^{-1}(\bm{u})$. Then
$V_{\mathrm{HK}}\left(h\right)\lesssim r^{d}\exp\left(M(1-\delta/2)^{2}r^{2}\right).$
Proof.
Let $\bm{x}=G^{-1}(\bm{u})$. For $\varnothing\neq\bm{v}\subseteq 1:d$, we have
$\partial^{\bm{v}}h(\bm{u})=\partial^{\bm{v}}(f\circ P_{r})(\bm{x})\prod_{j\in\bm{v}}\frac{dG_{j}^{-1}(u_{j})}{du_{j}}.$
By Lemma 3 of [basu2016], which is the special form of Faa di Bruno formula [cons1996], we have
$\partial^{\bm{v}}(f\circ P_{r})(\bm{x})=\sum_{\bm{\lambda}\in\mathbb{N}_{0}^{d},1\leq|\bm{\lambda}|\leq|\bm{v}|}D^{\bm{\lambda}}f(\bm{\vartheta})\sum_{s=1}^{|\bm{v}|}\sum_{(k_{l},\bm{\ell}_{l})\in{\mathrm{KL}}(s,\bm{v},\bm{\lambda})}\prod_{l=1}^{s}\partial^{\bm{\ell}_{l}}(P_{r}(\bm{x}))_{k_{l}},$
where $\bm{\vartheta}=P_{r}(\bm{x})$ and
${\mathrm{KL}}(s,\bm{v},\bm{\lambda})=\left\{\begin{array}[]{r}\left(\bm{\ell}_{l},k_{l}\right),l=1,\ldots,s,\mid\bm{\ell}_{l}\subseteq 1:d,k_{l}\in 1:d,\cup_{l=1}^{s}\bm{\ell}_{l}=\bm{v}\\ \quad\bm{\ell}_{l}\cap\bm{\ell}_{l^{\prime}}=\varnothing\text{ for }l\neq l^{\prime}\text{ and }\left|\left\{j\in 1:s\mid k_{j}=i\right\}\right|=\lambda_{i}\end{array}\right\}.$ (14)
By (ii) of Lemma 3.4, for any $\varnothing\neq\bm{v}\subseteq 1:d$ and $1\leq i\leq d$, we have $|\partial^{\bm{v}}(P_{r}(\bm{x}))_{i}|\lesssim\mathds{1}_{\{\|\bm{x}\|<r\}}.$ Hence,
$\left|\prod_{l=1}^{s}\partial^{\bm{\ell}_{l}}(P_{r}(\bm{x}))_{k_{l}}\right|\lesssim\mathds{1}_{\{\|\bm{x}\|<r\}}.$
By (iii) of Lemma 3.4, it follows that
$|D^{\bm{\lambda}}f(\bm{\vartheta})|\lesssim\exp(M\|P_{r}(\bm{x})\|^{2})\leq\exp\left(M(1-\delta/2)^{2}r^{2}\right).$
Therefore, we have
$|\partial^{\bm{v}}h(\bm{u})|\lesssim\exp\left(M(1-\delta/2)^{2}r^{2}\right)\mathds{1}_{\{\|\bm{x}\|<r\}}\prod_{j\in\bm{v}}\frac{1}{g_{j}(G_{j}^{-1}(u_{j}))}.$
Then, we consider $\int_{[0,1]^{|\bm{v}|}}|\partial^{\bm{v}}h\left(\bm{u}_{\bm{v}};\bm{1}_{\bar{\bm{v}}}\right)|d\bm{u}_{\bm{v}}$. If $\bar{\bm{v}}\neq\varnothing$, then $\bm{x}=G^{-1}(\bm{u}_{\bm{v}};\bm{1}_{\bar{\bm{v}}})=(\bm{x}_{\bm{v}},\bm{\infty}_{\bar{\bm{v}}})$, which means that $\|\bm{x}\|=\infty>r$ and $|\partial^{\bm{v}}h\left(\bm{u}_{\bm{v}};\bm{1}_{\bar{\bm{v}}}\right)|=0$. For $\bm{v}=1\!:\!d$, we have
$\int_{[0,1]^{|\bm{v}|}}\left|\partial^{\bm{v}}h\left(\bm{u}_{\bm{v}};\bm{1}_{\bar{\bm{v}}}\right)\right|d\bm{u}_{\bm{v}}=$ $\int_{[0,1]^{d}}\left|\partial^{1:d}h\left(\bm{u}\right)\right|d\bm{u}$
$\lesssim$ $~\exp\left(M(1-\delta/2)^{2}r^{2}\right)\int_{[0,1]^{d}}\mathds{1}_{\{\|\bm{x}\|<r\}}\prod_{j=1}^{d}\frac{1}{g_{j}(G_{j}^{-1}(u_{j}))}d\bm{u}$
$=$ $~\exp\left(M(1-\delta/2)^{2}r^{2}\right)\int_{\|\bm{x}\|<r}\prod_{j=1}^{d}\frac{1}{g_{j}(x_{j})}g_{j}(x_{j})d\bm{x}$
$\lesssim$ $~r^{d}\exp\left(M(1-\delta/2)^{2}r^{2}\right).$
In the last inequality, the spherical coordinate transformation was used. Hence, by the definition of the variation in the sense of Hardy and Krause [basu2016], we have
$V_{\mathrm{HK}}\left(h\right)$ $\leq\sum_{\varnothing\neq\bm{v}\subseteq 1:d}\int_{[{0},{1}]^{|\bm{v}|}}\left|\partial^{\bm{v}}h\left(\bm{u}_{\bm{v}};\bm{1}_{\bar{\bm{v}}}\right)\right|d\bm{u}_{\bm{v}}$
$=\int_{[0,1]^{d}}\left|\partial^{1:d}h\left(\bm{u}\right)\right|d\bm{u}\lesssim r^{d}\exp\left(M(1-\delta/2)^{2}r^{2}\right).$
This completes the proof. ∎
By K-H inequality defined in (4), we have the following Lemma.
Lemma 3.7.
Consider the setting of Lemma 3.6. Assume that $\left\{\bm{u}_{1},\ldots,\bm{u}_{N}\right\}$ is the RQMC point set such that the star discrepancy of $\{\bm{u}_{i}\}_{i=1}^{N}$ is almost surely $\mathcal{O}(N^{-1}(\log N)^{d-1})$, we have
$\mathbb{E}\left|{I}_{N}(f\circ P_{r})-\mathbb{E}[f\circ P_{r}(\bm{X})]\right|^{p}$ $\lesssim r^{pd}\exp\left(pM(1-\delta/2)^{2}r^{2}\right)\frac{(\log N)^{p(d-1)}}{N^{p}}.$
Now we turn to the proof of part (ii) of Theorem 1.
Proof of Theorem 1.
Combining the results of Lemmas 3.5 and 3.7, the error $\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}$ can be bounded by
$\mathcal{O}\left(r^{\eta}\exp\left(-pA(1-\delta)^{2}r^{2}\right)+r^{pd}\exp\left(pM(1-\delta/2)^{2}r^{2}\right)\frac{(\log N)^{p(d-1)}}{N^{p}}\right),$
where $A=\alpha/p-M>0$, and $(1-\delta)r>1/\sqrt{2pA}$. We focus on the two dominant exponential terms
$\exp\left(-pA(1-\delta)^{2}r^{2}\right)+\exp\left(pM(1-\delta/2)^{2}r^{2}\right)N^{-p}.$
Let $r=\sqrt{\theta\log N}$. The above equation is bounded by
$\mathcal{O}(N^{-pA(1-\delta)^{2}\theta}+N^{-p+pM(1-\delta/2)^{2}\theta}).$
To make the two terms comparable, we take
$\theta=\theta^{*}=\frac{1}{(1-\delta)^{2}A+(1-\delta/2)^{2}M},$
yielding an upper bound $\mathcal{O}(N^{-p\gamma_{1}})$, where
$\gamma_{1}=\frac{A}{A+\left(1+\frac{\delta}{2(1-\delta)}\right)^{2}M}.$
Hence, by setting $r=\sqrt{\theta^{*}\log N}$, it follows that
$\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}=\mathcal{O}\left({N^{-p\gamma_{1}}}(\log N)^{\max\{\eta/2,p(3d/2-1)\}}\right).$
Note that for any $\epsilon_{1}>0$, $(\log N)^{\max\{\eta/(2p),(3d/2-1)\}}=\mathcal{O}(N^{\epsilon_{1}})$. Moreover, since $\delta$ defined in (11) is sufficiently small, for any $\epsilon_{2}>0$, $\gamma_{1}=\frac{A}{A+M}-\epsilon_{2}=1-pM/\alpha-\epsilon_{2}$. Hence, for any $\epsilon>0$,
$\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left({N^{-\gamma+\epsilon}}\right).$
where $\gamma=1-pM/\alpha$. This completes the proof. ∎
For convenience, we unify the two cases of Theorem 1 into the following Corollary.
Corollary 1.
Given $p\geq 1$. Assume that $\mathcal{T}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$ and that $f$ satisfies Assumption 1 with $M<\alpha/p$. Then for any $\epsilon>0$, we have $\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$ where $\beta=1-p\max{\{M,0\}}/\alpha$. Moreover, if $M<\alpha/(2p)$, then $\beta>1/2$.
Next, we consider a general transport map $\mathcal{T}(\bm{u})=\tau(G^{-1}(\bm{u}))$ where $\tau:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a transformation and the inversion transformation $G^{-1}$ is defined in (7). Note that
$\pi(f)=\mathbb{E}_{\pi}\left[f(\bm{X})\right]=\mathbb{E}\left[f(\tau(G^{-1}(\bm{u}))\right]=\mathbb{E}\left[(f\circ\tau)(G^{-1}(\bm{u}))\right]:=\mathbb{E}\left[f_{\tau}(G^{-1}(\bm{u}))\right],$
where $f_{\tau}=f\circ\tau$. If $f_{\tau}$ satisfies Assumption 1 with $M\in\mathbb{R}$, then we can derive the $L_{p}$-error rate for general transport map $\mathcal{T}$ following Corollary 1.
Assumption 3.
Assume that for any $\bm{a}\in\mathbb{N}_{0}^{d}$ satisfying $1\leq|\bm{a}|\leq d$, there exists constants $C_{\tau},M_{\tau},C\in\mathbb{R}$ such that,
$\|\tau(\bm{x})\|\leq C_{\tau}\|\bm{x}\|+C,\quad|D^{\bm{a}}\tau_{i}(\bm{x})|\lesssim\exp(M_{\tau}\|\bm{x}\|^{2}),$
where $\tau_{i}$ denotes the $i$-th component of $\tau$.
Theorem 2.
Assume that $f$ satisfies Assumption 1 with $M_{f}\in\mathbb{R}$, and that $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$, then for any $\varepsilon>0$, $f_{\tau}=f\circ\tau$ satisfies Assumption 1 with a growth rate
$M_{f,\tau}=M_{f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}.$ (15)
Proof.
By Faa di Bruno formula [cons1996], for any $\bm{x}\in\mathbb{R}^{d}$ and $\bm{a}\in\mathbb{N}_{0}^{d}$, $D^{\bm{a}}f_{\tau}(\bm{x})$ is a sum of finite terms of the form
$D^{\bm{\lambda}}f(\tau(\bm{x}))\prod_{l=1}^{s}\prod_{i=1}^{d}(D^{\bm{\ell}_{l}}\tau_{i}(\bm{x}))^{\bm{k}_{l,i}},$ (16)
where $\bm{\lambda}\in\mathbb{N}_{0}^{d},1\leq|\bm{\lambda}|\leq|\bm{a}|$, $1\leq s\leq|\bm{a}|$, $\bm{k}_{l}\in\mathbb{N}_{0}^{d}\setminus\{{\bm{0}}\}$ satisfying $\sum_{l=1}^{s}\bm{k}_{l}=\bm{\lambda}$, and $\bm{\ell}_{l}\in\mathbb{N}_{0}^{d}\setminus\{{\bm{0}}\}$ such that $\sum_{l=1}^{{s}}|\bm{k}_{l}|\bm{\ell}_{l}=\bm{a}$. By Assumptions 3, it follows that
$\prod_{l=1}^{s}\prod_{i=1}^{d}\left|D^{\bm{\ell}_{l}}\tau_{i}(\bm{x})\right|^{\bm{k}_{l,i}}\lesssim\prod_{l=1}^{s}\exp(M_{\tau}|\bm{k}_{l}|\|\bm{x}\|^{2})=\exp(M_{\tau}|\bm{\lambda}|\|\bm{x}\|^{2})\leq\exp(dM_{\tau}\|\bm{x}\|^{2}).$
Hence, by Assumptions 1, we have
$|D^{\bm{a}}f_{\tau}(\bm{x})|\lesssim\exp\left(M_{f}\|\tau(\bm{x})\|^{2}\right)\exp\left(dM_{\tau}\|\bm{x}\|^{2}\right)\lesssim\exp\left((M_{f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau})\|\bm{x}\|^{2}\right).$
This completes the proof. ∎
Corollary 2.
Assume that the transport map $\mathcal{T}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$ and that transformation $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$. Suppose that $f$ satisfies Assumption 1 with growth rate $M_{f}$. If $M_{f,\tau}<\alpha/p$, then for any $\epsilon>0$, we have $\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$ where $\beta=1-p\max{\{M_{f,\tau},0\}}/\alpha$ and $M_{f,\tau}$ is defined in (15).
Now, we consider a specific form of the transformation $\tau(\bm{x})=T(L\bm{x}+\bm{\mu})$, where $T(\bm{x})=\left(T_{1}(x_{1}),T_{2}(x_{2}),\ldots,T_{d}(x_{d})\right)$ is a component-wise nonlinear transformation with each $T_{i}$ being a nonlinear function, $L\in\mathbb{R}^{d\times d}$ is a lower-triangular matrix (scaled transformation), and $\bm{\mu}\in\mathbb{R}^{d}$ (shifted transformation).
Proposition 1.
Assume that for $i=1,2,\ldots,d$, $T_{i}$ has $d$ times continuous derivatives and for any $\bm{v}\subseteq 1\!:\!d$, there exist constants $C_{T},M_{T},C_{1}\in\mathbb{R}$ such that,
$|T_{i}(y)|\leq C_{T}|y|+C_{1},\quad|D^{m}T_{i}(y)|\lesssim\exp(M_{T}|y|^{2}),\quad\forall y\in\mathbb{R},$ (17)
where $1\leq m\leq d$ and $1\leq i\leq d$. Then, for any $\varepsilon>0$, $\tau(\bm{x})=T(L\bm{x}+\bm{\mu})$ satisfies Assumption 3 with $C_{\tau}=C_{T}\sqrt{M_{L}}$ and $M_{\tau}=M_{T}M_{L}(1+\varepsilon)$, where $M_{L}=\lambda_{\max}(L^{\mathrm{T}}L)$.
Proof.
Let $L_{i}$ be the $i$-th row of $L$. Then, $\tau_{i}(\bm{x})=T_{i}(L_{i}\bm{x}+\mu_{i})$. By Faa di Bruno formula [cons1996], for any $\bm{x}\in\mathbb{R}^{d}$ and $\bm{a}\in\mathbb{N}_{0}^{d}$ satisfying $1\leq|\bm{a}|\leq d$, we have
$D^{\bm{a}}\tau_{i}(\bm{x})=D^{|\bm{a}|}T_{i}(y_{i})\prod_{j=1}^{d}(L_{i,j})^{a_{j}},$
where $y_{i}=L_{i}\bm{x}+\mu_{i}$. Since $|D^{m}T_{i}(y)|\lesssim\exp(M_{T}|y|^{2})$ for any $1\leq m\leq d$, it follows that for any $\varepsilon>0$,
$|D^{\bm{a}}\tau_{i}(\bm{x})|\lesssim\exp(M_{T}|L_{i}\bm{x}+\mu_{i}|^{2})\lesssim\exp(M_{T}M_{L}(1+\varepsilon)\|\bm{x}\|^{2}),$
where $M_{L}=\lambda_{\max}(L^{\mathrm{T}}L)$ and we use the fact that
$(L_{i}\bm{x})^{2}=\bm{x}^{\mathrm{T}}L_{i}^{\mathrm{T}}L_{i}\bm{x}\leq\lambda_{\max}(L_{i}^{\mathrm{T}}L_{i})\|\bm{x}\|^{2}\leq\lambda_{\max}(L^{\mathrm{T}}L)\|\bm{x}\|^{2}.$
In addition, since $|T_{i}(y)|\leq C_{T}|y|+C_{1},1\leq i\leq d$, we have
$\|\tau(\bm{x})\|\leq C_{T}\|L\bm{x}+\bm{\mu}\|+C_{1}\sqrt{d}\leq C_{T}\sqrt{M_{L}}\|\bm{x}\|+C_{T}\|\bm{\mu}\|+C_{1}\sqrt{d}.$
This completes the proof. ∎
Remark 5.
Consider the case of $T$ satisfying (17) and the standard Gaussian as the base distribution. Let $\bm{X}=\mathcal{T}(\bm{u})=T(\bm{y})$ where $\bm{y}=L\Phi^{-1}(\bm{u})+\bm{\mu}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ and $\bm{\Sigma}=L^{T}L$. For any $\varepsilon>0$, $\bm{X}$ satisfies Assumption 2 with $\eta=d-2$ and $\alpha=\lambda_{\min}(\bm{\Sigma}^{-1})/(2C_{T}^{2})-\varepsilon$. Indeed, note that $\|\bm{X}\|\leq C_{T}\|\bm{y}\|+\sqrt{d}C_{1}$. Then, for any $t>C_{T}/\sqrt{\lambda_{\min}(\bm{\Sigma}^{-1})}+\sqrt{d}C$ and $\varepsilon>0$, it follows that
$\mathbb{P}\left(\|\bm{X}\|>t\right)$ $\leq\mathbb{P}\left(C_{T}\|\bm{y}\|+\sqrt{d}C>t\right)=\mathbb{P}\left(\|\bm{y}\|>\frac{t-\sqrt{d}C}{C_{T}}\right)$
$\lesssim t^{d-2}\exp\left(-\left(\frac{\lambda_{\min}(\bm{\Sigma}^{-1})}{2C_{T}^{2}}-\varepsilon\right)t^{2}\right),$
where (8) was used in the last inequality.
Remark 6.
If $T(\bm{x})=\bm{x}$, i.e., $\tau(\bm{x})=L\bm{x}+\bm{\mu}$, then $C_{T}=1$ and $M_{T}=0$. Therefore, our results hold for all location–scale families that satisfy the sub-Gaussian tail Assumption 2, such as the classical Gaussian distribution. For such families, we have $C_{\tau}=\sqrt{M_{L}}$ and $M_{\tau}=0$. When the integrand $f$ satisfies Assumption  1 with $M_{f}\in\mathbb{R}$, then the composed function $f_{\tau}$ satisfies Assumption  1 with $M_{f,\tau}=M_{T}M_{L}(1+\varepsilon)$. Moreover, if $M_{f}<0$, then our results hold for all location–scale families.
Remark 7.
It is straightforward to check that the continuous nonlinear functions $T_{i}$ in Table 1 commonly used as the activation function in machine learning [sharma2017activation] satisfy (17).
Table 1: Common activation functions in machine learning.
Function Name Function $C_{T}$ $M_{T}$ $C_{\tau}$ $M_{\tau}$ $M_{f,\tau}$
Sigmoid $1/(1+\exp(-x))$ 0 0 0 0 0
Softplus $\log(1+\exp(x))$ 0 0 0 0 0
Tanh $\tanh(x)$ 1 0 $\sqrt{M_{L}}$ 0 $M_{f}M_{L}(1+\varepsilon)$
Swith $x/(1+\exp(-x))$ 1 0 $\sqrt{M_{L}}$ 0 $M_{f}M_{L}(1+\varepsilon)$
More general, we consider the transport map $\mathcal{T}=\tau(G^{-1}(\bm{u}))$ proposed by Liu [liu2024transport], which is used to approximate the target distribution. Here, $\tau=\tau^{1:K}=\tau^{K}\circ\cdots\circ\tau^{2}\circ\tau^{1}$ where $K\geq 1$, $\tau^{k}(\bm{x})=T^{k}(L^{k}\bm{x}+\bm{\mu}^{k})$, $T^{k}$ is a component-wise nonlinear transformation with $T^{k}(\bm{x})=(T_{1}^{k}(x_{1}),T_{2}^{k}(x_{2}),\ldots,T_{d}^{k}(x_{d}))$, $L^{k}\in\mathbb{R}^{d\times d}$ is a lower-triangular matrix and $\bm{\mu}^{k}\in\mathbb{R}^{d}$, $k=1,2,\ldots,K$.
Proposition 2.
If each $T^{k}_{i}$ satisfies (17) with the same parameters $C_{T}$ and $M_{T}$, then for any $\varepsilon>0$, $\tau=\tau^{K}\circ\cdots\circ\tau^{2}\circ\tau^{1}$ satisfies Assumption 3 with constant $C_{\tau}=(C_{T}\sqrt{M_{L}})^{K}$ and the correspond growth rate $M_{\tau}=M_{\tau}^{(K)}$ given recursively by
$M_{\tau}^{(K)}=M_{T}M_{L}(1+\varepsilon)C_{*}^{(K-1)}+dM_{\tau}^{(K-1)},~~M_{\tau}^{(0)}=0,~C_{*}^{(0)}=1,$
where $M_{L}=\max\limits_{1\leq k\leq K}\{\lambda_{\max}((L^{k})^{\mathrm{T}}L^{k})\}$, $C_{*}^{(K-1)}=(C_{T}\sqrt{M_{L}})^{2(K-1)}(1+\varepsilon)$.
Proof.
Similar to the proof of Theorem 4.5 of [liu2024transport], we establish the result by induction on $K$. By Proposition 1, we know that for $K=1$, the result holds. Assume that the result is true for K1K-1. For $i=1,2,\ldots,d$, define functions $H_{i}(\bm{x})=T_{i}^{K}(L_{i}^{K}\bm{x}+{\mu}_{i}^{K}),1\leq i\leq d$. By Faa di Bruno formula [cons1996], for any $\bm{x}\in\mathbb{R}^{d}$ and $\bm{a}\in\mathbb{N}_{0}^{d}$ satisfying $1\leq|\bm{a}|\leq d$, we have $D^{\bm{a}}\tau_{i}(\bm{x})=D^{\bm{a}}(H_{i}\circ\tau^{1:K-1})(\bm{x})$ can be expressed as a finite sum of terms of the form
$(D^{\bm{\lambda}}H_{i}(\tilde{\bm{x}}))\prod_{l=1}^{s}\prod_{i=1}^{d}(D^{\bm{\ell}_{l}}\tau^{1:K-1}_{i}(\bm{x}))^{\bm{k}_{l,i}},$ (18)
where $\tilde{\bm{x}}=\tau^{1:K-1}(\bm{x})$, $\bm{\lambda}\in\mathbb{N}_{0}^{d},1\leq|\bm{\lambda}|\leq|\bm{a}|$, $1\leq s\leq|\bm{a}|$, $\bm{k}_{l}\in\mathbb{N}_{0}^{d}\setminus\{{\bm{0}}\}$ satisfying $\sum_{l=1}^{s}\bm{k}_{l}=\bm{\lambda}$, and $\bm{\ell}_{l}\in\mathbb{N}_{0}^{d}\setminus\{{\bm{0}}\}$ such that $\sum_{l=1}^{{s}}|\bm{k}_{l}|\bm{\ell}_{l}=\bm{a}$. By the induction assumption on $\tau^{1:K-1}$, it follows that $\|\tilde{\bm{x}}\|=\|\tau^{1:K-1}(\bm{x})\|\lesssim(C_{T}\sqrt{M_{L}})^{K-1}\|\bm{x}\|+C$ and
$\prod_{l=1}^{s}\prod_{i=1}^{d}\left|D^{\bm{\ell}_{l}}\tau^{1:K-1}_{i}(\bm{x})\right|^{\bm{k}_{l,i}}\lesssim\exp\left(dM_{\tau}^{(K-1)}\|\bm{x}\|^{2}\right).$ (19)
Again, by Faa di Bruno formula [cons1996], we have
$D^{\bm{\lambda}}H_{i}(\tilde{\bm{x}})=D^{|\bm{\lambda}|}T_{i}^{K}(y_{i})\prod_{j=1}^{d}(L_{i,j})^{\lambda_{j}},$
where $y_{i}=L_{i}^{K}\tilde{\bm{x}}+{\mu}_{i}^{K}$, Hence,
[…]
$\lesssim\exp\left(M_{T}M_{L}(1+\varepsilon)\left(\left(C_{T}\sqrt{M_{L}}\right)^{2(K-1)}(1+\varepsilon)\right)\|\bm{x}\|^{2}\right).$
By (18) and (19) together with (20), we have
$|D^{\bm{a}}\tau_{i}(\bm{x})|\lesssim\exp(M_{T}M_{L}(1+\varepsilon)C_{*}^{(K-1)}\|\bm{x}\|^{2})\exp\left(dM_{\tau}^{(K-1)}\|\bm{x}\|^{2}\right)=\exp\left(M_{\tau}^{(K)}\|\bm{x}\|^{2}\right).$
Moreover, since $|T_{i}^{K}(y)|\leq C_{T}|y|+C_{1},1\leq i\leq d$, it follows that
$\|\tau(\bm{x})\|$ $=\|T^{K}(L^{K}\tilde{\bm{x}}+\bm{\mu}^{K})\|\leq C_{T}\|L^{K}\tilde{\bm{x}}+\bm{\mu}^{K}\|+\sqrt{d}C_{1}$
$\leq C_{T}\sqrt{M_{L}}\|\tilde{\bm{x}}\|+C_{T}\|\bm{\mu}^{K}\|+\sqrt{d}C_{1}$
$\lesssim C_{T}\sqrt{M_{L}}(C_{T}\sqrt{M_{L}})^{(K-1)}\|\bm{x}\|+C_{T}\sqrt{M_{L}}C+C_{T}\|\bm{\mu}^{K}\|+\sqrt{d}C_{1}.$
This completes the proof. ∎
For $\tau=\tau^{K}\circ\cdots\circ\tau^{2}\circ\tau^{1}$, by Proposition 2, we see that $M_{\tau}$ is recursively calculated. But as shown in Remark 7, for commonly used activation functions, $M_{T}=0$. Hence, $M_{\tau}$ remains zero for all $K$.
By Proposition 2 and Remark 7, our theoretical framework applies to neural network mappings constructed from commonly used activation functions, thereby providing a solid mathematical foundation for extending the theoretical advantages of RQMC to various deep learning models.
4 The $L_{p}$-error rate of RQMC-based SNIS estimator
In this section, we study the $L_{p}$-error rate of RQMC-SNIS estimator for unbounded function $f$. We consider the case of target distribution $\pi=\bar{\pi}/Z$ with an unknown normalizing constant $Z>0$. Given a proposal $q$, define $\omega=\bar{\pi}/q$. We rewrite $\pi(f)$ as
$\pi(f)=\int_{\mathbb{R}^{d}}\pi(\bm{x})f(\bm{x})d{\bm{x}}=\frac{1}{Z}\int_{\mathbb{R}^{d}}\bar{\pi}(\bm{x})f(\bm{x})d{\bm{x}}=\frac{1}{Z}\int_{\mathbb{R}^{d}}\omega(\bm{x})f(\bm{x})q(\bm{x})d{\bm{x}}:=\frac{q(\omega f)}{q(\omega)},$
where $q(\omega)=\int_{\mathbb{R}^{d}}\omega(\bm{x})q(\bm{x})d\bm{x}=\int_{\mathbb{R}^{d}}\bar{\pi}(\bm{x})d{\bm{x}}=Z$ and $q(\omega f)=\int_{\mathbb{R}^{d}}\omega(\bm{x})f(\bm{x})q(\bm{x})d\bm{x}$. We also rewrite the RQMC-SNIS estimator $\pi_{N}(f)$ as
$\pi_{N}(f)=\sum_{i=1}^{N}\frac{\omega\left(\bm{\bm{x}}_{i}\right)}{\sum_{j=1}^{N}\omega\left(\bm{\bm{x}}_{j}\right)}f\left(\bm{\bm{x}}_{i}\right)=\frac{q_{N}(\omega f)}{q_{N}(\omega)},$
where $q_{N}(\omega)=\frac{1}{N}\sum_{i=1}^{N}\omega\left(\bm{\bm{x}}_{i}\right)$, $\bm{x}_{i}=\mathcal{T}(\bm{u}_{i})\sim q$, $\left\{\bm{u}_{1},\ldots,\bm{u}_{N}\right\}$ denotes the RQMC point set satisfying (6) and each $\bm{u}_{i}\sim\mathcal{U}(0,1)^{d}$, and transform map $\mathcal{T}=\tau\circ G^{-1}$ with transformation $\tau$ and $G^{-1}$ defined in (7).
Before we give the $L_{p}$-error rate of RQMC-SNIS, we need to establish the $L_{p}$-error rate for RQMC integration with functions $\omega$ and $\omega f$.
Lemma 4.1.
Assume that $1/q$, $\bar{\pi}$, and $f$ each satisfy Assumption 1 with growth rates $M_{q},~M_{\pi}$, and $M_{f}\in\mathbb{R}$, respectively. Then $\omega$ and $\omega f$ each satisfy Assumption 1 with respective growth rates $M_{\omega}=M_{q}+M_{\pi}$ and $M_{\omega f}=M_{q}+M_{\pi}+M_{f}$.
Proof.
By the Leibnitz’s formula [cons1996], for any $\bm{a}\in\mathbb{N}_{0}^{d}$, it follows that
$|D^{\bm{a}}\omega(\bm{x})|$ $=\sum_{\bm{0}\leq\bm{b}\leq\bm{a}}\binom{\bm{a}}{\bm{b}}|D^{\bm{b}}\bar{\pi}(\bm{x})||D^{\bm{a-b}}(1/q)(\bm{x}))|$
$\lesssim\sum_{\bm{0}\leq\bm{b}\leq\bm{a}}\binom{\bm{a}}{\bm{b}}\exp\left(M_{\pi}\|\bm{x}\|^{2}\right)\exp\left(M_{q}\|\bm{x}\|^{2}\right)$
$\lesssim\exp\left((M_{q}+M_{\pi})\|\bm{x}\|^{2}\right),$
where $\bm{0}\leq\bm{b}\leq\bm{a}$ means $0\leq b_{i}\leq a_{i}$ for $i=1,2,\ldots,d$. Similarly, we have
$|D^{\bm{a}}(\omega f)(\bm{x})|\lesssim\exp\left((M_{q}+M_{\pi}+M_{f})\|\bm{x}\|^{2}\right).$
This completes the proof. ∎
By Corollary 1 and Theorem 2, we have the following result.
Corollary 3.
Assume that the transport map $\mathcal{T}=\tau\circ G^{-1}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$, and that transformation $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$. Suppose that $1/q$, $\bar{\pi}$, and $f$ satisfy Assumption 1 with respective growth rates $M_{q},M_{\pi}\in\mathbb{R}$, and $M_{f}>0$. If $M_{\omega f,\tau}<\alpha/p$, then for any $\epsilon>0$, we have
$\left\{\mathbb{E}\left|q_{N}(\omega f)-q(\omega f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$
where $\beta=1-p\max{\{M_{\omega f,\tau},0\}}/\alpha$ and $M_{\omega f,\tau}$ is defined in (15).
If the unknown constant $q(\omega)=1$ and $\pi=\bar{\pi}$, then $q(\omega f)=\pi(f)$. Corollary 3 can be also seen as the $L_{p}$ error rate for RQMC based IS estimator.
Remark 8.
In Section 5 of Ouyang et al. [ouyang2024], they also employed their projection method to derive the error bound of RQMC-IS. For the target distribution $\pi\sim\mathcal{N}(\bm{0},I_{d})$, they assumed that $\mathcal{T}=G^{-1}(u)\sim q$, and that for any $\varnothing\neq\bm{v}\subseteq 1\!:\!d$, $|\partial^{\bm{v}}(f/q)(\bm{x})|\lesssim\exp(M|\bm{x}|^{2})$ where $M<1/2$, which implies $|\partial^{\bm{v}}(\omega f)(\bm{x})|\lesssim\exp\big((M-1/2)\|\bm{x}\|^{2}\big).$ Under these assumptions, they derived that the $L_{2}$ error for RQMC-IS is $\mathcal{O}\big(N^{-1}(\log N)^{3d/2-1}\big)$. Clearly, their result corresponds exactly to the case in Corollary 3 with $M_{\omega f,\tau}=M_{\omega f}<0$ and $p=2$.
By Lemma 4.1, Gaussian-type proposals will increase the overall growth rate of the integrand, thereby may reducing the achievable convergence rate. He et al. [he2024] considered a proposal generated by a linear transformation of IID t-distributed random variables with degrees of freedom $\nu>0$. Specifically, for a location vector $\bm{\mu}$ and a positive definite matrix $\bm{\Sigma}$, the transport map $\mathcal{T}$ is defined as $\mathcal{T}(\bm{u})=\bm{\mu}+LF_{t}^{-1}(\bm{u}),$ where $LL^{T}=\bm{\Sigma}$ and the components of $F_{t}^{-1}$ denote the inverse CDF of the uni-variate t-distribution. For such a proposal $q$, $C_{\tau}=\sqrt{M_{L}}$ and $M_{\tau}=0$. Moreover, $1/q$ satisfies Assumption 1 with arbitrarily small $M_{q}>0$. For the importance sampling case, assume that $\pi(\bm{x})=\bar{\pi}(\bm{x})$ satisfies Assumption 1 with growth rate $M_{\pi}<0$. Consequently, if unbounded functions $f$ satisfy Assumption 1 with growth rate $0<M_{f}<-M_{\pi}$, then we obtain $M_{\omega f}=M_{f}+M_{\pi}+M_{q}<0$ and $M_{\omega f,\tau}<0$. Hence, by Corollary 3 for any $\epsilon>0$, $\left\{\mathbb{E}\left|q_{N}(\omega f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right)$ . In particular, for Gaussian target distributions $\pi(\bm{x})\sim N(\bm{\mu},\bm{\Sigma})$, we have $M_{\pi}=-\frac{1}{2}\lambda_{\min}(\bm{\Sigma}^{-1})+\varepsilon<0$. The condition $0<M_{f}<-M_{\pi}$ and the rate $\mathcal{O}\left(N^{-1+\epsilon}\right)$ are consistent with Theorem 5.1 in [he2024].
The main result of this work is following. For convenience, we focus on the case of $M_{\omega,\tau}>0$ is arbitrarily small and $M_{f}>0$.
Theorem 3.
Consider the setting of Corollary 3. Assume that $M_{\omega,\tau}>0$ is arbitrarily small and $M_{f}>0$. Suppose that there exists a constant $v>1$ such that $q\left(|f|^{v}\right)=\int_{\mathbb{R}^{d}}|f(\bm{x})|^{v}q(\bm{x})d\bm{x}<\infty$. If $1\leq p<\min\left\{v,\alpha/M_{\omega f,\tau}\right\}$, then for any $\epsilon>0$, the RQMC-SNIS estimator has a $L_{p}$-error rate
$\left\{\mathbb{E}\left|\pi_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$
where $\beta=1-pM_{\omega f,\tau}/\alpha>0.$ Moreover, if $M_{\omega f,\tau}<\alpha/(2p)$, then $\beta>1/2$.
Proof.
To obtain the result of Theorem 3, we follow the proof strategy of [deligiannidis2024, Theorem 2.3]. Define
$\mathcal{E}:=\left\{\left|q_{N}(\omega)-q(\omega)\right|>0.5\min(1,q(\omega))\right\}.$
By the law of total probability,
$\mathbb{E}\left[\left|\pi_{N}(f)-\pi(f)\right|^{p}\right]=\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}}\right]+\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}^{c}}\right].$ (21)
For the first term, we have
$\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}}\right]\leq$ $~\mathbb{E}\left[\left(\max_{1\leq i\leq n}f\left(\bm{x}_{i}\right)+\frac{q(\omega f)}{q(\omega)}\right)^{p}\mathds{1}_{\mathcal{E}}\right]$
$\leq$ $~2^{p-1}\mathbb{E}\left[\left(\max_{1\leq i\leq n}f\left(\bm{x}_{i}\right)\right)^{p}\mathds{1}_{\mathcal{E}}\right]+2^{p-1}\left(\frac{q(\omega f)}{q(\omega)}\right)^{p}\mathbb{P}(\mathcal{E}).$
By Corollary 3 and Markov inequality, for any constant $c>0$ and $1\leq s<\frac{\alpha}{M_{\omega,\tau}}$, we have
$P\left(\left|q_{N}(\omega)-q(\omega)\right|>c\right)\leq\frac{\mathbb{E}\left[\left|q_{N}(\omega)-q(\omega)\right|^{s}\right]}{c^{s}}=\mathcal{O}\left(N^{-(1-\epsilon)s}\right).$
Hence, $\mathbb{P}(\mathcal{E})=\mathcal{O}\left(N^{-(1-\epsilon)s}\right).$ Then, by Holder’s inequality with exponents $v/p$ and $(1-p/v)^{-1}$, we have
$\mathbb{E}\left[\left(\max\limits_{1\leq i\leq N}f\left(\bm{x}_{i}\right)\right)^{p}\mathds{1}_{\mathcal{E}}\right]\leq$ $~\mathbb{E}\left[\left(\max\limits_{1\leq i\leq N}|f\left(\bm{x}_{i}\right)|\right)^{v}\right]^{p/v}\times\left[\mathbb{P}(\mathcal{E})\right]^{1-p/v}$
$\leq$ $~q\left(|f|^{v}\right)^{p/v}\mathcal{O}\left(N^{p/v-(1-\epsilon)s(1-p/v)}\right),$
where we use the fact that
$\mathbb{E}\left[\left(\max\limits_{1\leq i\leq N}|f\left(\bm{x}_{i}\right)|\right)^{v}\right]\leq\mathbb{E}\left[\sum_{i=1}^{N}\left|f\left(\bm{x}_{i}\right)\right|^{v}\right]=Nq(|f|^{v}).$
To ensure
$p/v-(1-\epsilon)s(1-p/v)\leq-(1-\epsilon)p,$ (22)
it is required that $1\leq p\leq\dfrac{sv}{s+v+1/(1-\epsilon)}$. Note that $\frac{sv}{s+v+1/(1-\epsilon)}$ is monotonically increasing for $s>0$ and
$\lim\limits_{s\to\infty}\frac{sv}{s+v+1/(1-\epsilon)}=v.$
Since $M_{\omega,\tau}>0$ is arbitrarily small, we take a sufficiently large $s$ such that (22) holds for any $1\leq p<v$. It then follows (22) that for any $s\geq 1$,
$(1-\epsilon)s\leq-(1-\epsilon)p.$
Hence,
$\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}}\right]=\mathcal{O}\left(N^{-(1-\epsilon)p}\right).$ (23)
For the second term of (21), note that
$\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|\leq\frac{1}{|q_{N}(\omega)|}\left|q_{N}(f\omega)-q(f\omega)\right|+\left|\frac{q(f\omega)}{q_{N}(\omega)q(\omega)}\right|\left|q_{N}(\omega)-q(\omega)\right|,$
and
$\mathcal{E}^{c}$ $=\left\{\left|q_{N}(\omega)-q(\omega)\right|<0.5\min(1,q(\omega))\right\}$
$=\left\{q(\omega)-0.5\min(1,q(\omega))<q_{N}(\omega)<q(\omega)+0.5\min(1,q(\omega))\right\}.$
Therefore, for any $\epsilon>0$,
$\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}^{c}}\right]\lesssim$ $~\mathbb{E}\left[\left|q_{N}(f\omega)-q(f\omega)\right|^{p}\right]+\mathbb{E}\left[\left|q_{N}(\omega)-q(\omega)\right|^{p}\right]$ (24)
$=$ $~\mathcal{O}\left(N^{-(\beta-\epsilon)p}\right)+\mathcal{O}\left(N^{-(1-\epsilon)p}\right)=\mathcal{O}\left(N^{-(\beta-\epsilon)p}\right),$
where $\beta=1-\frac{M_{\omega f,\tau}}{\alpha/p}<1$ since $p<\frac{\alpha}{M_{\omega f,\tau}}$. By (21) and (23) together with (24), the result follows immediately. ∎
Remark 9.
An unbounded function $f$ that satisfies Assumption 2 with an arbitrarily small growth rate $M>0$ is referred to as a “QMC-friendly” function [he2024]. If function $\omega f$ is QMC-friendly, then $M_{\omega f}>0$ is arbitrarily small. When $M_{\tau}=0$, $M_{\omega f,\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)$ remains arbitrarily small. Since $M_{\omega f}=M_{\omega}+M_{f}$ with $M_{f}>0$, we have $M_{\omega}<M_{\omega f}$ and $M_{\omega,\tau}<M_{\omega f,\tau}$, and thus $M_{\omega,\tau}$ is arbitrarily small. By Theorem 3, for any $\epsilon>0$ and $1\leq p<v$, we have $\left\{\mathbb{E}\left|\pi_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right).$ Compared with the $L_{p}$-error rate $\mathcal{O}(N^{-1/2})$ in Theorem 2.3 of [deligiannidis2024], our result achieves a faster rate. Moreover, we require weaker moment condition on $f$ and $\omega$. In Theorem 2.3 of [deligiannidis2024], for constant $s\geq 2$, it needs $q(\omega^{s})<\infty$ and the condition on $p$ is $2\leq p\leq\frac{sv}{s+v+2}$. For instance, consider the case of $p=2$ (RMSE) and $s=3$, [deligiannidis2024] needs $v=10$ to achieve an error rate of $\mathcal{O}(N^{-1/2})$. In contrast, if function $\omega f$ is QMC-friendly, our result only requires $v=2+\delta$ to achieve an error rate of $\mathcal{O}(N^{-1+\epsilon})$ for any $\delta>0$ and $\epsilon>0$.
5 Numerical experiments
In this section, we investigate a toy Bayesian inverse problem and a Bayesian logistic regression problem to examine the effect of the parameters in Theorem 3 for RQMC-SNIS, driven by classical scrambling Sobol’ point sets with direction numbers from [joe2003], which satisfies (6). We repeat the estimator independently $R$ times to estimate the $L_{p}$-error rate of the estimators. The empirical $L_{p}$ error is computed as
$\widehat{L_{p}}=\left[\frac{1}{R}\sum_{i=1}^{R}\left|\pi^{(i)}_{N}(f)-\hat{\pi}(f)\right|^{p}\right]^{1/p},$ (25)
where $\pi^{(i)}_{N}(f)$ are IID copies of the RQMC estimator $\pi_{N}(f)$, and
$\hat{\pi}(f)=\frac{1}{R}\sum_{i=1}^{R}\pi^{(i)}_{N}(f)$
denotes the final estimate of the posterior expectation $\pi(f)$. In the following experiments, we fix $R=50$, $f(\bm{x})=\|\bm{x}\|^{2}$ which satisfies Assumption 1 with arbitrarily small $M_{f}>0$, and we consider a linear transform map $\mathcal{T}(\bm{u})=\bm{\mu}+LG^{-1}(\bm{u})$ with $\bm{\mu}\in\mathbb{R}^{d}$ and $L\in\mathbb{R}^{d\times d}$ satisfying $L^{\mathrm{T}}L=\bm{\Sigma}$ for a positive definite matrix $\bm{\Sigma}$. Hence, for any $\varepsilon>0$, $M_{\omega f,\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}=M_{\omega f}M_{L}(1+\varepsilon)$ where $M_{L}=\lambda_{\max}(\bm{\Sigma})$.
5.1 Bayesian inverse problem
Consider a Bayesian inverse problem with the forward model $\bm{y}=\mathcal{G}(\bm{z})+\bm{\varsigma}$, where $\bm{y}\in\mathbb{R}^{d}$ is the observed data, $\bm{z}\in\mathbb{R}^{d}$ is the unknown parameter, $\mathcal{G}$ is the response operator, and $\bm{\varsigma}\in\mathbb{R}^{d}$ is the noise vector. Assume that $\bm{\varsigma}\sim\mathcal{N}(\bm{0},(1/n)I_{d})$, where $n$ is the level of noise. The likelihood $l(\bm{z})$ is proportion to $\exp\left\{-n\Psi(\bm{z})\right\}$ with $\Psi(\bm{z})=\frac{1}{2}\|\bm{y}-\mathcal{G}(\bm{z})\|^{2}$. Let $\pi_{0}(\bm{z})$ be prior distribution. By Bayes’ formula, the posterior distribution is given by
$\pi(\bm{z})=\frac{\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})}{\int_{\mathbb{R}^{d}}\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}},$
where the denominator is usually an intractable constant.
Our aim is to estimate the posterior expectation of interest given by
$\pi(f):=\mathbb{E}_{\pi}[f(\bm{Z})]=\frac{\int_{\mathbb{R}^{d}}f(\bm{z})\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}{\int_{\mathbb{R}^{d}}\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}.$
Following [he2024], we consider a Gaussian prior $\pi_{0}=N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$ and a linear mapping with a small nonlinear perturbation as the response operator to analyze the effect of the model parameters. Specifically, we take $\mathcal{G}(\bm{z})=\bm{z}+\lambda\mathcal{F}(\bm{z})$ with $\lambda>0$ and $\mathcal{F}(\bm{z})=\left(z_{1}e^{-z_{1}^{2}},\ldots,z_{d}e^{-z_{d}^{2}}\right)^{\mathrm{T}}$. For simplicity, we take $\bm{y}=0$, yielding
$\Psi(\bm{z})=\frac{1}{2}\|\mathcal{G}(\bm{z})\|^{2}=\frac{1}{2}\sum_{i=1}^{d}z_{i}^{2}\left(1+\lambda e^{-z_{i}^{2}}\right)^{2}.$
Note that $\Psi(\bm{z})\geq\frac{1}{2}\|\bm{z}\|^{2}$. By Example 4.9 of [he2024], $\Psi(\bm{z})$ satisfies Assumption 1 with arbitrarily small $M_{\Psi}>0$. Hence, the likelihood $l(\bm{z})$ satisfies Assumption 1 with $M_{l}=-\frac{n}{2}+M_{\Psi}<0$. For the Gaussian prior $N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$, we set $\bm{\mu_{0}}=\bm{1}$ and $\bm{\Sigma}_{0}=I_{d}$. Therefore, the posterior $\pi(\bm{z})$ has a growth rate $M_{\pi}=-\frac{1}{2}-\frac{n}{2}+M_{\Psi}<0$.
Note that $\bm{\mu}^{*}=\mathbf{0}$ is the global minimizer of $\Psi(\bm{z})$, and $\bm{\Sigma}^{*}=\left(\nabla^{2}\Psi\left(\bm{\mu}^{*}\right)\right)^{-1}=\kappa I_{d}$ with $\kappa=(1+\lambda)^{-2}$. In our experiments, we vary the model parameter $\kappa$, the dimension $d$, and the moment parameter $p$ to show the performance of various RQMC-SNIS estimators. We consider five proposals as follows:
  • PriorIS: Gaussian proposal, $\bm{\mu}=\bm{\mu_{0}}=\mathbf{1}$ and $\bm{\Sigma}=\bm{\Sigma}_{0}=I_{d}$;
  • ODIS: Gaussian proposal, $\bm{\mu}=\bm{\mu}^{*}=\mathbf{0}$ and $\bm{\Sigma}=\bm{\Sigma}_{0}=I_{d}$;
  • LapIS: Gaussian proposal, $\bm{\mu}=\bm{\mu}^{*}=\mathbf{0}$ and $\bm{\Sigma}=\bm{\Sigma}^{*}/n=(\kappa/n)I_{d}$;
  • tIS0: Linear t proposal, $\bm{\mu}=\mathbf{1}$, $\bm{\Sigma}=I_{d}$ and the degrees of freedom $v=5$;
  • tIS: Linear t proposal, $\bm{\mu}=\mathbf{0}$, $\bm{\Sigma}=(\kappa/n)I_{d}$ and the degrees of freedom $v=5$.
For simplicity, throughout the experiments, we fix $n=20$. More impact with different noise $n$, we refer to [he2024].
For LapIS, we have $M_{L}=\kappa/n$ and $M_{q}=\frac{n}{2\kappa}$. Applying Lemma 4.1, then we have
$M^{\mathrm{LapIS}}_{\omega f}=M_{f}+\frac{n}{2\kappa}-\frac{1}{2}-\frac{n}{2}+M_{\Psi}=\frac{n}{2\kappa}\left(1-\frac{n+1}{n}\kappa\right)+\varepsilon_{1},$
where $\varepsilon_{1}=M_{f}+M_{\Psi}>0$ is arbitrarily small. If $\kappa>\frac{n}{n+1}$, then, $M^{\mathrm{LapIS}}_{\omega f,\tau}=M^{\mathrm{LapIS}}_{\omega f}M_{L}\leq 0$. Otherwise, $M^{\mathrm{LapIS}}_{\omega f,\tau}>0$ which is not arbitrarily small. For tIS and tIS0, $M_{q}>0$ is arbitrarily small. Then we have
$M^{\mathrm{tIS0}}_{\omega f}=M^{\mathrm{tIS}}_{\omega f}=M_{f}+M_{q}+M_{\pi}=M_{f}+M_{q}-\frac{1}{2}-\frac{n}{2}+M_{\Psi}<0.$
Hence, for any $p$ and $\kappa$, $M^{\mathrm{tIS0}}_{\omega f,\tau}=M^{\mathrm{tIS}}_{\omega f,\tau}<0$. For PriorIS and ODIS, we also have $M^{\mathrm{PriorIS}}_{\omega f,\tau}=M^{\mathrm{ODIS}}_{\omega f,\tau}<0$ for any $p$ and $\kappa$. Therefore, by Theorem 3, for LapIS proposal with $\kappa>\frac{n}{n+1}$ and other four proposals, we can get $\mathcal{O}(N^{-1+\epsilon})$ rate for $L_{p}$-error.
We first consider the case $d=5$, $\kappa=1/4$ and $p\in\{1,2,5\}$. Figure 2(a) reports $\log_{2}\widehat{L_{p}}$ given by (25) with $\log_{2}N$. For the RQMC-SNIS with proposals PriorIS, ODIS, and tIS0, we observe an empirical error decay of order $\mathcal{O}(N^{-1/2})$, similar to standard MC. This deviates from the theoretical predictions, primarily because the associated variances are excessively large, which leads to implicit constants in the error rate that cannot be neglected and thus causing deviations from the expected rates.
Figure 2(b) illustrates the corresponding effective sample sizes. It is clear that PriorIS, ODIS, and tIS0 result in very small effective sample sizes, whereas only tIS and LapIS achieve effective sample sizes comparable to the actual sample sizes. From 2(a), we see that LapIS performs poorly when $\kappa$ is small, and its performance deteriorates as $p$ increases; in contrast, tIS consistently achieves the expected $\mathcal{O}(N^{-1+\epsilon})$ convergence rate across different $p$, matching the theoretical predictions.
In the following, since the parameter choices for PriorIS, ODIS, and tIS0 are independent of $\kappa$, we focus only on LapIS and tIS for $\kappa\in\{3/4,1\}$. We observe that LapIS with $\kappa=3/4$ outperforms the previous case with $\kappa=1/4$, though its performance still deteriorates as $p$ increases. When $\kappa=1$ satisfying $\kappa>\frac{n}{n+1}$, the results of LapIS become independent of $p$, consistently exhibiting a stable convergence rate of $\mathcal{O}(N^{-1+\epsilon})$. The convergence rate of tIS remains stable and unaffected by both $\kappa$ and $p$, in alignment with theoretical expectations.
(a) $\kappa=1/4$.
(b) Effective sample sizes with $\kappa=1/4$.
(c) $\kappa=3/4$.
(d) $\kappa=1$.
Figure 2: The $L_{p}$-error for different proposals with different $\kappa$ and $d=5$.
Finally, we consider the case $d=30$ to assess the effect of dimensionality in Figure 3, focusing on $\kappa=1$. As the dimension increases, the convergence rates of both LapIS and tIS inevitably deteriorate due to the $(\log N)^{(3d/2-1)}$ factor in the rate. While this factor can be absorbed into $N^{\epsilon}$ for sufficiently large $N$, it still has a noticeable impact when $N$ is relatively small. Developing an RQMC–SNIS algorithm whose performance is dimension-independent remains an open problem. Recently, Pan et al. [pan2025quasi] proposed a boundary-damping importance sampling approach for QMC integration, focusing on component-wise target distributions. Their method can achieve an RMSE rate of order $\mathcal{O}(N^{-1+\epsilon})$ that are indepedent of dimension. This work provides a promising direction for addressing the dimensionality challenge in RQMC–SNIS.
Figure 3: The $L_{p}$-error for different proposals with $\kappa=1$ and $d=30$.
5.2 Bayesian logistic regression
We consider a Bayesian logistic regression model following [he2023]. Let $\bm{X}\in\mathbb{R}^{m\times d}$ denote the observation matrix of predictor variables with $\bm{X}_{i}\in\mathbb{R}^{1\times d}$ the $i$-th row, $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{d})^{T}\in\mathbb{R}^{d\times 1}$ denote the regression parameters, and $\bm{Y}=(Y_{1},Y_{2},\ldots,Y_{m})$ denote the observation with $Y_{i}\in\{0,1\}$. The logistic model is expressed as
$\mathbb{P}\left(Y_{i}=1\right)=1-\mathbb{P}\left(Y_{i}=0\right)=\frac{\exp(\bm{X}_{i}\bm{\beta})}{1+\exp(\bm{X}_{i}\bm{\beta})},\ i=1,2,\ldots,m.$
The log-likelihood function for $\bm{Y}$ is given by
$\log l(\bm{\beta})=\bm{\beta}^{T}\bm{X}^{T}\bm{Y}-\sum_{i=1}^{m}\log\left(1+\exp\left(\bm{X}_{i}\bm{\beta}\right)\right).$
We take a standard Gaussian prior $\bm{\beta}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$. We consider the Pima Indian dataset with $m=392,d=9$ [windle2013]. Following [he2023], we compare two proposal distributions:
  • LapIS: A Gaussian proposal with mean $\bm{\mu}_{\star}$ and covariance $\bm{\Sigma}_{\star}$, where $\bm{\mu}_{\star}$ solves
    $\bm{\mu}_{\star}=\sum_{i=1}^{m}\left(Y_{i}-\frac{\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}{1+\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}\right)\bm{X}_{i}^{T},$
    and the covariance $\bm{\Sigma}_{\star}$ is given by
    $\bm{\Sigma}_{\star}=\left(I_{d}-\nabla^{2}F\left(\bm{\mu}_{\star}\right)\right)^{-1},$
    where $F(\bm{\beta})=\log l(\bm{\beta})$ and
    $\nabla^{2}F\left(\bm{\mu}_{\star}\right)_{jk}=-\sum_{i=1}^{m}\frac{\bm{X}_{ij}\bm{X}_{ik}\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}{\left(1+\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)\right)^{2}}.$
  • Linear t proposal, $\bm{\mu}=\bm{\mu}_{\star},\bm{\Sigma}=\bm{\Sigma}_{\star}$.
By Faa di bruno formula [cons1996], we can check that the likelihood function $l(\bm{\beta})$ satisfies Assumption 1 with arbitrarily small $M_{l}>0$. Hence, $M_{\pi}=-1/2+M_{l}$. Then, for any $\varepsilon_{1}>0$ and $\varepsilon_{2}>0$, we have
$\alpha^{\mathrm{LapIS}}=\lambda_{\min}(\bm{\Sigma}_{\star}^{-1})/2-\varepsilon_{1},\quad M^{\mathrm{LapIS}}_{\omega f,\tau}=\frac{1}{2}\left(\lambda_{\max}(\bm{\Sigma}_{\star}^{-1})-1\right)\lambda_{\max}(\bm{\Sigma}_{\star})(1+\varepsilon_{2})>0,$
where we have use the fact that $\lambda_{\max}(\bm{\Sigma}_{\star}^{-1})>1$ by Theorem 5.3 of [he2023]. However, we cannot guarantee that $\alpha^{\mathrm{LapIS}}/M^{\mathrm{LapIS}}_{\omega f,\tau}>1$, and consequently our theoretical results do not apply to LapIS in this model. On the other hand, for tIS we have $M^{\mathrm{tIS}}_{\omega f,\tau}=-\frac{1}{2}\lambda_{\max}(\bm{\Sigma}_{\star})+\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0$. Taking a small enough $\varepsilon^{\prime}$ such that $M^{\mathrm{tIS}}_{\omega f,\tau}<0$, the tIS proposal achieves an $L_{p}$-error rate of nearly $\mathcal{O}(N^{-1})$. Figure 4 illustrates the estimate of $L_{p}$-error for $p=\{1,2,5\}$. It can be observed that LapIS performs worse than tIS, whereas the convergence rates of tIS remain consistently close to $\mathcal{O}(N^{-1})$ across different values of $p$.
Figure 4: The $L_{p}$-error for different proposals with Pima dataset.
6 Conclusions
In this work, we establish $L_{p}$-error rate $(p\geq 1)$ of RQMC-SNIS estimator in the setting of unbounded integrands on $\mathbb{R}^{d}$ for general test distributions. We validate our theoretical results through a simple Bayesian inverse problem, examining the impact of various parameters, and find that the experimental outcomes align well with the theory. In addition, we conduct experiments on Bayesian logistic regression, where the results suggest that the linear $t$ proposals constructed using $\bm{\mu}$ and $\bm{\Sigma}$ determined by the LapIS method, appear to provide superior performance. Our results fill an important theoretical gap in the foundation of RQMC-SNIS and broaden its applicability to practical Bayesian computation.
In the course of this derivation, we also establish $L_{p}$-error rate for RQMC integration for a broad class of transport maps, including shifted, scaled, and commonly used nonlinear transformations, as well as to a wide class of test functions satisfying the critical quadratic growth. This advancement broadens the theoretical scope of RQMC methods beyond the standard normal setting. Moreover, the established $L_{p}$-error bounds provide the analytical basis for higher-order moment analysis and uncertainty quantification, offering a rigorous foundation for future applications of RQMC techniques. More applications of $L_{p}$-error for RQMC see Appendix A.
However, in the experiments, we observed that SNIS estimators constructed directly with low-discrepancy point sets still perform poorly in high-dimensional problems. Developing an RQMC-SNIS algorithm that is robust to dimensionality remains an important direction for future research.
Appendix A Some applications of $L_{p}$-error for RQMC
Using the $L_{p}$-error for RQMC integration, we can bound the skewness $\gamma$ and the kurtosis $\kappa$ in RQMC, which are important quantities in some properties of confidence intervals. For convenience, we focus on the case $f$ is “QMC-friendly”, which means that for any $\epsilon>0$ and $p\geq 1$, $\left\{\mathbb{E}\left|{I}_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right).$ Assume that $\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]=\Omega(N^{-2})$. Then
$|\gamma|=\frac{|\mathbb{E}[({I}_{N}(f)-\pi(f))^{3}]|}{\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]^{3/2}}\leq\frac{\mathbb{E}[|{I}_{N}(f)-\pi(f)|^{3}]}{\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]^{3/2}}=\frac{\mathcal{O}\left(N^{-3+\epsilon}\right)}{\Omega\left(N^{-3}\right)}=\mathcal{O}\left(N^{\epsilon}\right),$ (26)
and,
$\kappa=\frac{\mathbb{E}[({I}_{N}(f)-\pi(f))^{4}]}{\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]^{2}}=\frac{\mathcal{O}\left(N^{-4+\epsilon}\right)}{\Omega\left(N^{-4}\right)}=\mathcal{O}\left(N^{\epsilon}\right).$ (27)
Based on these, we can further analyze some properties of confidence intervals for RQMC. Recently, it has been observed that a standard Student’s t based confidence interval is empirically more effective than bootstrap t interval [l2023confidence]. Pan and Owen [pan2025skewness] explained the empirical phenomenon by analyzing the skewness of RQMC. For a review of confidence intervals in the RQMC setting, see [owen2024error].
In RQMC, we independently repeat the estimator $R$ times, yielding $R$ IID random variables $X_{i}={I}_{N}^{(i)}(f)$. For $0<a<1/2$, the standard Student’s t based confidence interval for $\pi(f)$ with nominal coverage 12a1-2a takes the form
$\bar{X}\pm t_{(R-1)}^{1-a}S/\sqrt{R},$
where
$\bar{X}=\frac{1}{R}\sum_{i=1}^{R}X_{i},\quad S^{2}=\frac{1}{R-1}\sum_{i=1}^{R}\left(X_{i}-\bar{X}\right)^{2},$
and $t_{(R-1)}^{1-a}$ is the 1a1-a quantile of the Student’s $t$ distribution on R1R-1 degrees of freedom.
Pan and Owen [pan2025skewness] considered the skewness for RQMC directly without an absolute value in the numerator. As a consequence, the result of [pan2025skewness] cannot be applied to the classical Berry–Esseen bound [bentkus1996berry]. In contrast, our estimation of skewness is based on the absolute third moment, which ensures that the Berry–Esseen bound can be applied. Define the Student’s t statistic $T_{R}=\frac{\bar{X}-\pi(f)}{S/\sqrt{R}}$. A Berry-Esseen type bound for $T_{R}$ quantifies the worst-case deviation of its distribution to the standard normal distribution, given by $\Gamma_{R}=\sup\limits_{x\in\mathbb{R}}|\mathbb{P}\left(T_{R}<x\right)-\Phi(x)|.$ By Theorem 1.1 of [bentkus1996berry], there exists an absolute constant $c>0$ such that
$\sup\limits_{x\in\mathbb{R}}\left|\mathbb{P}\left(\frac{\bar{X}-\pi(f)}{\hat{\sigma}/\sqrt{R}}<x\right)-\Phi(x)\right|\leq\frac{c}{\sqrt{R}}\frac{\mathbb{E}[|{I}_{N}(f)-\pi(f)|^{3}]}{\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]^{3/2}},$
where $\hat{\sigma}^{2}=\frac{R-1}{R}S^{2}$. Hence,
$\Gamma_{R}$ $=\sup\limits_{x}\left|\mathbb{P}\left(\frac{\bar{X}-\pi(f)}{\hat{\sigma}/\sqrt{R}}<\sqrt{\frac{R-1}{R}}x\right)-\Phi(x)\right|$
$\leq\sup\limits_{x}\left|\mathbb{P}\left(\frac{\bar{X}-\pi(f)}{\hat{\sigma}/\sqrt{R}}<\sqrt{\frac{R-1}{R}}x\right)-\Phi\left(\sqrt{\frac{R-1}{R}}x\right)\right|+\left|\Phi\left(\sqrt{\frac{R-1}{R}}x\right)-\Phi(x)\right|$
$\leq\frac{c}{\sqrt{R}}\frac{\mathbb{E}[|{I}_{N}(f)-\pi(f)|^{3}]}{\mathbb{E}[({I}_{N}(f)-\pi(f))^{2}]^{3/2}}+\mathcal{O}(R^{-1})=\mathcal{O}(N^{\epsilon})\mathcal{O}(R^{-1/2})+\mathcal{O}(R^{-1}).$
Since $\epsilon>0$ is arbitrarily small, $\mathcal{O}(N^{\epsilon})$ is almost $\mathcal{O}(1)$. Then, $\Gamma_{R}\approx\mathcal{O}(R^{-1/2})$.
Another important quantity is the coverage error in the standard confidence interval, defined as
$\kappa_{R}=\mathbb{P}\left(\bar{X}-\frac{S}{\sqrt{R}}t_{(R-1)}^{1-a}\leqslant\pi(f)\leqslant\bar{X}+\frac{S}{\sqrt{R}}t_{(R-1)}^{1-a}\right)-(1-2a).$
By [owen2025coverage], the coverage error in Student’s t interval is given by
$\frac{z_{1-a}}{R}\left(\frac{z_{a}^{2}-3}{6}\kappa-\frac{z_{a}^{4}+2z_{a}^{2}-3}{9}\gamma^{2}\right)\Phi^{\prime}\left(z_{a}\right)+\mathcal{O}\left(R^{-3/2}\right).$
where $z_{a}=\Phi^{-1}(a)$ and $\Phi^{\prime}(x)$ is the standard normal density. By (26) and (27), we have $\kappa_{R}=\mathcal{O}(N^{2\epsilon})\mathcal{O}(R^{-1})+\mathcal{O}(R^{-3/2}).$ Since $\epsilon$ can be arbitrarily small, the convergence rate of $\kappa_{R}$ is nearly $\mathcal{O}(R^{-1/2})$.